A Complete Bibliography of Publications in Mathematical Tables and Other Aids to Computation, 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

    E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
               WWW URL: https://www.math.utah.edu/~beebe/

17 September 2024 Version 2.28

Title word cross-reference

(−ϵ/T) [CFH58]. (1 + x^{2})^{1/2} [Ano57h]. $0.20. [Goo57b]. $1.00 [Mel57]. $1.35 [Hen57a]. $10.00 [Dav57]. $12.50 [Hen57b]. 15 [BMSW54, HS55]. 16 [BMSW54, Lin57]. 2, 000 [MRvN50]. $2.00 [T.57e]. 25 [Sus59]. 25 × 34 [Tom57]. 27 × 48 [Wil57b]. 2^{1/3} [vNT55]. 2^{n} ± 1 [Rob57]. 2 ln(1 + x^{2}) [Ano57h]. $5.00 [T.57b]. $5.86 [Gre57b]. $6.50 [F.57, F.57]. $7 [F.57]. $8.00 [T.57a]. 812 × 11 [Hol57b]. 81 × 11 [Wil57b]. 9n [Por56]. arcsin z [Ano57d]. arc tan Z [Luk57]. arctg [Ano57h]. AX^{2} − I = 0 [Laa58]. d^{2} [GM51]. ẍ = f(x, y, z, t) [Her51]. e [MRvN50, Rei50a]. e^{z} [Ive53]. e_{m}(S_{n}) [Wyn56].

Ei(x) [Har57a]. ∫_{0}^{∞} e^{−x}J_{0}(ηxξ)J_{1}(xξ)x^{−n}dx

[Wee59]. ∫_{0}^{∞} e^{−x}x^{n}f(x)dx [RW59]. ∫_{x}^{∞} J_{0}(t)t^{−n}dt [Lon59]. ∫_{x}^{∞} J_{1}(t)t^{−n}dt [Lon59]. ln Γ[z] [Ano57f]. ln x [Ano57h]. log Z [Luk57]. n^{2} + 1 [Sha59b]. π [MRvN50, NJ55, Rei50a]. ψ(θ) = ∑_{n=1}^{∞} sin nθn^{2} [AS56]. s [Tha57]. ℘ [Sou57, T.57f]. x/1 − x^{2} [GCN59]. x^{2} [GCN59]. x^{3} + y^{3} + z^{3} = 1 [New57a].

. [Dia57]. .C [Boo54].

1-3 [New57c]. 160 [New57c]. 163 [T.57g].

1867-corrected [Hen57b]. 1953. [Ano57f]. 1957 [Ano57e].

233-250 [New57c]. 247 [Ano57d]. 280 [New57a]. 298 [New57b].

454 [T57d]. 485 [Coh57].

50th [Gau94]. 55 [Ewa57b]. 59-72 [New57c].

617 [Ewa57a]. 650 [Wag57].

701 [Joh56]. 709 [Wal58].

A. [Ano57f, F.57, T.57c]. A.E.C.L [Mel57]. A.P.E [Boo54]. Abh [T.57h]. Abramov [Ano57f]. Abschatzungen [New57c]. Abscissas [RW59]. Access [Per52]. Accuracy [Kun56, Lot51, Sal51c, SF59, Ste53, Was58]. Accurate [Lon57]. Acht [Ano57d]. Acht- [Ano57d]. Acta [New57b]. Adams [ZS59]. Addendum [BY53]. Addition [Mon50]. Adjoint [Laa59]. Advanced [Est53]. Adventures [BGW52]. Aids [AL51, Gep52, MW51, Max54, Mon50, SR53]. Airy [RB59]. Aitken [Twe54]. Akad [Ano57f, Rho57, T.57h]. Akademiia [Ano57h, Ano57g]. Alan [Tau56]. Albert [Hen57a]. Algebraic [Cur54, Gep52, Lin57, Mul56, Pos55, She58, de 53]. Algebraic-Hyperbolic [Lin57]. Algorithm [Sha54]. All-Purpose [D. 53]. Alternating [CD58]. Analog [Max54, Neu59]. Analogue [SR53]. Analogues [PN58]. Analyser [Atk55]. Analysis [Ano50h, BWW54, CO51, Fox51, FM51, Goo57b, HMW55, Hus50, Tau50]. Analyst [For53]. Analytic [DR54, Gau59]. Analyzer [Don52, Kor52, Spr52]. angew [New57c]. Annals [Ano57d]. Anniversary [Gau94]. Anwendungen [New57c]. Application [Men57a]. Applications [Fle56a]. Applied [Goo57b, T.57e]. Appreciation [Gib50]. Approximate [Mor58, de 53]. Approximating [HH55, MW59, Boo55]. Approximation [Gro54, Har57b, Hol57a, Sou57, T.57f]. Approximations [BD59, Cle54, Hit57, She58, Was58]. Arbitrary [Gre55a]. Archibald [Ben50]. Argument [Sal51b, Sou57, T.57f]. Arguments [FM51, GT58, Kun56]. Arising [Luk56, Sal55, Wor55]. Army [Dia57]. Arsham [Dia57]. Associated [Sus59]. Asynchronous [Lub52]. Atomic [Mel57, Ski59]. August [Gau94]. Australia [Pea52]. Automatic [Ano50h, BGW52, BW53, Cah54, CO51, Dav57, Fox51, Har50, Lub52, MDL50, MS51, Mul56, Pea52, Per52, Sti50, Tod50, Wag57, WR50, Ano51f, Boo54, Coh50, Cou50, Don52, Ele50, Est53, FM51, Gar54, Gre53b, Hum54, Hus50, T.51, KN51, MDL50, Pea52, Spr52, Ste51, Sto53, Wal53].

B [Est57, T57d, Ano57d]. Back [Ano53a]. BARK [KN51]. Based [F.57, Men57b]. Bateman [F.57, Hen57a]. Battin [Dav57]. Bayer [T.57h]. Be [Sti50, SZF57]. beispielweise [Ewa57b]. Belford [Hol57b]. Belgique [T.57b]. Bennett [Hen57a]. benötigt [Ewa57b]. Ber [New57c]. Bessel [GT58, GT59, Hit57, Lon57, Luk59, SK58, SA57]. Between [Mor58]. Bi [Coh58]. Bi-Quadratic [Coh58]. Bibliography [Ano50a, Ano50b, Ano51a, Ano53b, Ano53c, Ano53d, Ano54a, Ano54b]. Bierens [Hen57b]. Biharmonic [CD58]. Binary [Har50]. binders [Hol57b]. Bivariate [Gau59, Sou56]. Blaise [Cou50]. Blunder [BGW52]. Body [GRS53]. Book [Dav57, Hen59, T.57c]. Books [Ano55d]. Boolean [Pos55]. Boundaries [Vis57]. Boundary [GL56, Hen59, Mit59]. Bounds [de 53]. British [Est57, Gau94]. Bureau [Tom57].

C [Dia57, Goo57b, Sch54, T57d, T.57e]. Calcul [T.57b]. calculated [MRvN50]. Calculating [LaF54, Sal51a, SS59]. Calculation [Fet55, GT59, MS51, MH58, Opl51, Pol59, SR53, SF59]. Calculations [Arc50]. Calculator [F. 50, Har50, MS51, Wag54, WR50]. California [Abb51, Mor51]. Canada [Mel57]. Canonical [Hea57]. capacity [CD53]. Card [Gar54, GS52, MS51, Opl51, Wag54]. Card-Programmed [Wag54]. Cards [BY51, BY53]. Carlo [DR56, Edm53, FL50, Opl51]. Cartesian [T.57g, Ano57h]. Case [Rei50b, Sou57, T.57f]. Cauchy [Lon58]. Cauer [T.57h]. Cauer-Parameter [T.57h]. Central [Sal59]. century [Gau94]. Certain [Bol57, Fet55, GS52, Lon57, Mor56, Sch59, SS59, WW57]. Chalk [Mel57]. Characteristic [SB53]. Characteristics [Ele50, Hus50]. Characters [BMSW54]. Chebyshev [Dia57, Chi56, Cle55, Dia57, MW59]. Checking [Mil50]. Chem [T57d]. Chemie [Ewa57b]. chisel [Ano57h]. Christoffel [Rei50b]. Circle [Gre53b]. Clare [Ben50]. Class [Coh58]. cm [Dav57, F.57, Goo57b, Gre57b, Hen57a, Hen57b, Mel57, Rho57, T.57c, T.57a, T.57b, T.57e, Tom57, Mel57]. cm. [Tom57]. Co [Dav57, Hen57b]. Codes [HMW55]. Coding [Ano53b, Ano54a]. Coefficient [Doy57]. Coefficients [Bla52, CFH58, Dia57, SS59, T.57e, T.57g]. Collector [Gre55b]. Columbia [Gau94]. Comment [Coh57, Dav57, Est57, Ewa57a, Ewa57b, F.57, Goo57b, Gre57b, Hen57a, Hen57b, Hol57b, Mel57, New57a, New57b, New57c, Rho57, T.57h, T57d, T.57f, T.57a, T.57b, T.57e, T.57g, Tom57, Wil57b]. Comments [NJ55]. Complete [MDL50]. Complex [Ano57f, Atk55, Der59, Men57a, Sal51a, Ano57h, T.57g]. Compressible [Ber55]. Computation [Ano57d, Coh58, Luk57, LWF72, Men57a, NJ55, SZF57, Spr52, Tom57, CD53, Gau94, AL51, Gep52, MW51, Max54, Mon50, SR53]. computational [Gau94]. Computations [Fro58b, Gar54]. Computer [Abb51, Est53, Est57, Gre53b, Hus50, KN51, MDL50, Mor51, Mul56, Neu59, Pea52, Wal53, Wol52, Sto53]. Computers [D. 53, Lub52, Max54, Mul58, SA57, Sti50, T57d]. Computing [Abb51, Ano50h, Ano51f, BR53, Boo54, BGW52, BW53, Cah54, CO51, Coh50, Cou50, DR56, Don52, Ele50, Est53, Fox51, FM51, Gar54, Gre53a, Gre53b, Har50, Hea57, Hum54, Hus50, T.51, KN51, Lub52, MDL50, MS51, Mor56, Neu59, Pea52, Per52, Rho57, Spr52, Ste51, Sti50, Sto53, Tei52, Tod50, Wal53, WW57, Wyn56]. Concepts [Spr52]. Concerning [Arc50, KR59, Goo55]. Condition [FM52, Tau50]. Conditioned [Ril55]. Conditions [DF53, Mit59]. Cones [HMS56]. Configurations [Coh56]. Confluent [Luk59, Mil57]. Conjecture [KR59, vNG53]. Conjectures [Goo55]. Conjugate [Cur54]. Considerations [F. 50]. Constant [Bla52]. Constants [Fis57]. Continued [Tei52, vNT55]. Control [Dav57, Dav57]. Convention [Est57]. Convergence [Fra50]. Conversion [Har50, Mel57, Sod53]. Convex [GJ59]. Convolution [Mir59]. coordinates [Ano57h]. copies [Hol57b]. Corps [Dia57]. corrected [Hen57b]. Corrector [Her51]. Correlations [Hea57]. Corrigenda [Ano50c, Ano50d, Ano50e, Ano50f, Ano51b, Ano51c, Ano51d, Ano51e, Ano52a, Ano52b, Ano52c, Ano53e, Ano53f, Ano53g, Ano53h, Ano54c, Ano54d, Ano55a, Ano55b, Ano55c, Ano56a, Ano56b, Ano56c, Ano57d, Ano57a, Ano57b, Ano57c, Ano58, Ano59a, Ano59b].

Corrigendum [Ano54e, Ano54f]. Couffignal [Gre57b]. Coupon [Gre55b]. Crank [T.57a]. Critical [Mit59]. cube [CD53]. Cumulative [LWF72]. Curve [Hol57a]. Curved [Vis57]. Cyclic [Ost53].

D [Dia57, Goo57b, T.57e]. D.

[Hen57b, New57a, New57b]. Decimal

[Har50, MS51, Rei50a, MRvN50]. Define [CFH58]. Définies [Hen57b]. Definite [Doy57, Fet55, Har57b, Ril55, Rit50, Wag54]. Degree [BMSW54]. Degrees [War57]. dekartovykh [Ano57h]. Delay [WR50]. Density [Lee58]. Department [Dia57]. deposited [Hol57b, Wil57b]. d’Équations [Gre57b]. Derivative [Sal51b]. Derivatives [CFH58, Ho55]. Descriptions [Ano55d]. Desiatiéna [Ano57h]. Desk [Har50]. Determinants [Por56]. Determination [Ber55, HS55, Laa59, MW59, Rei50a]. Development [Boo54]. Developments [Ano51f, Boo54, Coh50, Cou50, Don52, Est53, Gar54, Gre53b, Hum54, Hus50, KN51, MDL50, Pea52, Spr52, Ste51, Sto53, Wal53]. Deviates [Mul58, Tom56]. Device [Ost58, Wyn56]. Devices [BR53]. Diagonalization [Fro58a]. Diamond [Dia57]. Difference [Fox53, Kor52, Mit59, Ost52, Pol59, Ril54, She55, Was58]. Differences [Bla54a, Bla54b, Ho55, Kun56, Mil50, SK58, Sou56]. Differentiable [MW59]. Differential [Atk55, Bla52, Cam57, CO51, Don52, DF53, F. 50, FGT57, Fra50, HH55, Hen57a, Hen59, Kor52, Laa59, Spr52, Ste53, SM58, Tod50, Wil57a, Wil58, Wor55]. Differentiating [AL51]. Differentiation [SK58]. Diffusion [Wor55, T.57a]. Digital [Abb51, Coh50, Don52, Est57, Lub52, Mor51, Spr52, Wal53, T57d]. Digits [Gre55b, Tom56, Gru50, GM51, MRvN50]. Dimensional [BMv^{+}59]. Dimensions [Tha57]. d’Intégrales [Hen57b]. diophantine [New57a]. Direction [CD58]. Dirichlet [Ril54]. discriminant [Coh57]. Discussions [Ano50h, BGW52, BW53, Cah54, CO51, Fox51, Har50, Lub52, MS51, Per52, Sti50, Tod50]. Disposition [Str57]. Distribution [Coo57, Har57b, Sal59, Sha59a]. Divided [Bla54a, Bla54b, Kun56]. doppelperiodische [New57c]. Double [Ho55]. Dover [Hen57a]. Drum [Coh50]. Dynamic [BD59].

E. [T.57g]. Earth [Tom57]. Eastern [MDL50]. edition [Hen57b]. Editorial [T.51]. EDSAC [BW53, WR50]. Effect [Tom57]. Efficient [Ski59]. Eigensolutions [Laa59]. Eigenvalues [Gre53a, Neu59]. Eigenvectors [Gre53a, Neu59]. Eight [Ano57d, HSW56]. Electrical [Est57]. Electron [Mel57]. Electronic [Atk55, Est53, MS51, Rho57, WR50]. Electrons [Coh56]. Elektronnye [Rho57]. Elementary [Cle54]. Elliptic [She55]. Elliptical [Ano57d]. elliptischen [Ano57d]. Emersleben [Ewa57a, Ewa57b, New57c]. Energy [Mel57]. Eng [T57d]. Engineering [Dav57]. Engineers [Est57, Est57]. ENIAC [MRvN50, Rei50a]. Enumeration [Isb59]. Epsteinsche [New57c]. Epsteinscher [New57c]. Equally [Sal57]. Equally-Weighted [Sal57]. Equation [BLS57, CD58, Kor52, Laa58, Mil57, SB53, She58, Ski59, Vis57, Was58, Wor55, New57a, Hol57b]. Equations [BR53, Ber55, Bla52, Cam57, CO51, Cur54, DF53, F. 50, FGT57, FM52, Fox53, Fra50, Goo57a, GS52, HH55, Laa59, MW51, Men57b, Mul56, Pos55, Ril55, She55, Sny55, Ste53, SM58, Tod50, Wil57a, Wil58, de 53, Hen57a, Hen59]. Equianharmonic [Sou57, T.57f]. Equipped [Gar54]. Erd’elyi [F.57]. Ernst [T.57h]. Errata [Ano53i, Ano54g, Ano57f, Ano57h, Ano57g, Dia57]. Error [F. 50, Sal51a, ZS59, de 53]. Errors [DR54]. Especially [Gar54]. Estimates [Hoc58]. Estimation [DR54, Gar56]. Evaluating [Fle56b]. Evaluation [Bol57, Der59, Gre57a, HW57, HS58, Lon57, Lon58, Luk56, Mir59, Pos55, RW59, Sal55]. Events [Edm53]. Everett [Chi56, Sou56]. Excitation [Pol59]. Expansion [Fle56a, Luk59, Ste51, vNT55]. Expansions [SR53]. Experiments [DR56, Hoc58]. Explicit [Sus59]. Exponent [Por56]. Exponential [BM52a, BM52b, CFH58, FM51, Har57a, MH58]. Extension [FM52, SK58]. Extrapolation [ZS59].

F [F.57]. F. [T57d]. Factor [Gar56]. Factoring [Sha59b]. Factorizations [Rob57]. Factors [Sel53]. Fairchild [Sto53]. Fedorova [T.57c]. Fermat [Fro58b, Sel53]. Ferut [Hum54]. Field [Der59]. fields [Coh57]. Fifteen [For53]. Fifteen-Foot [For53]. Figure [Mel57]. File [Wil57b]. file. [Hol57b]. Finding [Atk55, Gre55a, SB53, Sal51b, Wil55]. Finite [Fox53, Ho55, Mit59, Pol59, Smi59]. Finite-Difference [Fox53]. First [Wil57a, Wil58, MRvN50]. Five [Ano57d, Mel57]. Flames [Wor55]. Floating [BW53, MS51]. Flow [Was58]. Fluids [Ber55]. Flüssigkeiten [New57c]. Foot [For53]. Form [DOH54, Doy57, MH57, RW59, Rob57, Sha59b]. Formula [Chi56, Fet55, Har57b, Rei50b, Sou56, Wil57a, Wil58]. Formulae [Coo57, Ho55, Sus59]. Formulas [Fie58, GCN59, Kun56, Sal51a, Sal51b, Sal56, Sal57, Str57, Tha57]. Fourier [HZ56, HPZ59]. Fourth [Sou56]. Fox [Hen59]. Fr.b [T.57b]. Fraction [vNT55]. Fractions [Tei52]. Free [BWW54]. Freedom [War57]. Frenelia [Ano57g]. Fresnel [Ano57g]. Function [Ano57d, AS56, Gre57a, GJ59, Har57b, Luk59, MW59, Sal51a, Sal51b, Sou57, T.57f, New57b]. Functional [BD59]. Functions [Ano57d, Cle54, DR54, Fle56a, Gau59, GT58, GT59, GCN59, Hit57, LaF54, Lon57, Luk59, Mor56, RB59, Sal59, Sal51c, SF59, SA57, Sus59, WW57, Ano57h, Boo55, Tom57, F.57]. Fundamental [Ber55, Spr52]. Fünfstellige [Ano57d]. Funktionen [Ano57d]. funktsiı [Ano57h]. Future [Goo57b].

G [F.57]. G. [Hol57b]. Games [Isb59]. Gamma [SF59]. Gauss [FM52, Leh56, Mel57, Ost58]. Gauss-Cm [Mel57]. Gaussian [RB59]. Gauthier [Gre57b]. Gauthier-Villars [Gre57b]. Gebelein [Ano57d]. Gellman [Mel57]. Genauigkeit [Ewa57b]. General [Dia57, KN51, Mil57]. Générale [T.57b]. Géneralisiées [T.57b]. Generalization [Cur54, Smi57]. Generalized [Smi59]. Generating [Joh56]. Generation [Mul58, SA57]. Given [Sal51b]. Glowatzki [T.57h]. Godwin [Coh57]. Gov [Goo57b]. Govt [T.57e]. Gradient [Vis57]. Gradients [Cur54]. Graphic [MW51]. Graphical [Mir59]. Greifswald [Ewa57a]. grid [T.57g]. Grids [Wil57b]. Group [PN58]. Groups [BMSW54, Lee58]. Guide [BY51, BY53, T.57c].

H [Ano57d, Dav57, F.57, Mel57, New57a, New57b, T57d, T.57f]. Haan [Hen57b]. Hafner [Hen57b]. Halcomb [Dav57]. half [Gau94]. half-century [Gau94]. Harry [Hen57a]. Harvard [Ano57d]. Having [Lee58]. Heat [Was58]. Hermitian [Fro58a]. Heuristic [PU59]. High [BR53, Kun56, Sal51c, SF59, SA57, Tei52, VR51]. Higher [F.57]. Hilbert [Smi59]. Hill [Dav57, F.57, Hol57b, BLS57]. Hollander [T57d]. Hulls [Pol59]. Hydrodynamical [BMv^{+}59]. Hydrodynamics [PU59]. Hyperbolic [Lin57]. Hypergeometric [Cah54, Luk59, Mil57].

IBM [Joh56, MS51, Wag57, Wal58]. II [Bla54b, BM52b, HS58, HPZ59, Tau50, Wil58]. iii [Gre57b]. Ill [Ril55]. Ill-Conditioned [Ril55]. ILLIAC [Gre53a, Sny55]. Implicit [Was58]. Improvement [Sny55]. Improving [FM52]. inches [Wil57b]. Incorporation [MDL50]. Indeterminacy [Cam57]. Index [LWF72]. Infinite [Bol57, Lon57]. Influence [Mit59]. Information [Ano53m, Arc50, Coh50, Rho57]. Informatksi [Rho57]. ington [T.57e]. Input [Hum54]. Institut [Ano57h, Ano57g, Cou50]. Institute [Est53, Hus50]. Institution [Est57]. Integral [FM51, Goo57a, GS52, Har57a, Luk56, MH58, Mir59, Rit50]. Intégrales [T.57b]. Integralov [Ano57g]. Integrals [Ano57g, Bol57, Dal54, DR56, Fet55, HW57, HS58, Har57b, HZ56, HPZ59, Lin57, Lon57, Lon58, Mor58, RW59, RB59, Sal57]. Integrating [Cam57]. Integration [Bar59, BM52a, BM52b, Fis57, GL56, GCN59, HMS56, HS56, Hen57a, Her51, Luk56, Pei57, SZF57, SM58, Str57, Wee59, Wil57a, Wil58]. Interfaces [She58]. Interpolating [Sal59]. Interpolation [Chi56, Gau59, Ho55, Ost52, Sou56, Twe54, T.57g, T.57e]. Interval [Sad50]. Intervals [Fox53]. Inverse [DOH54, LaF54, Mor56, Mul58, Sal55, WW57]. Inverses [Sal51c, Smi59, Wil57b]. Inversion [BR53, Doy57, Edm53, FL50, Sal57, Sch59, Was52]. Inverting [MH57, Wag54]. Investigate [Dij57]. Involving [Bla52, Lon57, RB59]. Irregular [Por56]. Isaac [Arc50]. Isabelle [Dia57]. issue [T.57f]. Iteration [Ost53, Ost58, Ril54]. Iterative [Fle56a, Fle56b, Fra50, Laa58, Men57b, Twe54, Wil55]. iv [Goo57b]. ix [Dav57, T.57e].

J [Hol57b, T.57g]. J. [Coh57, Dav57, Mel57, T.57a]. Jackson [Hol57b]. Jn [Coh57, New57a]. John [Ano57e]. Jr [Dav57, Wil57b]. July [Mel57]. Jury [War57].

Kind [Sus59]. Kinetic [Mel57]. Kl [T.57h]. know [T57d]. Kompleknoı [Ano57f]. kompleksnykh [Ano57h]. konstante [Ewa57b]. Konstanten [Ewa57a]. koordinat [Ano57h]. kopeks [Rho57]. Kristalls [Ewa57b]. Kummer [vNG53]. Kutta [Lot51].

L [Hen59, Hol57b, Mel57, Wil57b]. L. [Gre57b]. Laboratories [Dia57]. Laboratory [Ano57d]. Laethem [T.57b]. Lagrangian [T.57e]. Laning [Dav57]. Laplace [Sal55, She58]. Laplacian [Wil57b]. Large [FM51, Fox53, GT58, MH57, Mul58]. Large-Scale [Mul58]. Laslett [Hol57b]. Latent [Wil55]. leaves [Wil57b]. Lebedev [Rho57, T.57c]. left [F.57]. Legendre [Sus59]. Lehmer [New57a, New57b]. Lemma [Wag57]. Life [Arc50]. Like [Gro54]. Limit [Sal59]. Lindow [SK58]. line [Ano57d]. Linéaires [Gre57b]. Linear [BR53, Cur54, Dia57, FM52, Gau59, Gep52, MW51, Men57b, Ost52, Ril55, Sny55, SS59, Twe54, de 53]. Liouville [FGT57]. Location [Leh56]. logarifmov [Ano57h]. Logarithm [Sha54]. Logarithmic [Luk56, Mon50]. logarithms [Ano57h]. Logical [Abb51, BWW54]. London [Coh57, Est57, F.57, New57a]. Lotkin [Ano57d]. Louvain [T.57b]. Ltd [Mel57].

M [Ano57d, T.57c]. Machine

[AL51, BWW54, Cou50, Gar54]. Machinery [Ano50h, BGW52, BW53, Cah54, CO51, Fox51, Har50, Lub52, MS51, Per52, Sti50, Tod50, Ano51f, Boo54, Coh50, Cou50, Don52, Ele50, Est53, FM51, Gar54, Gre53b, Hum54, Hus50, T.51, KN51, MDL50, Pea52, Spr52, Ste51, Sto53, Wal53]. Machines [Opl51, Rho57]. Madelung [Ewa57a, Ewa57b]. Madelung-konstante [Ewa57b]. Magnetic [Coh50]. Magnus [F.57]. Mahalanobis [Sch54]. Majority [Isb59]. Making [FM51]. man [Ewa57b]. MANIAC [HMW55]. Many [War57]. Marcel [T.57b]. Mark [Ano57d]. Mashiny [Rho57]. Matematicheskie [Ano57h, Ano57g]. Matematicheskim [T.57c]. Math [Coh57, Ewa57a, New57a, New57c, T.57h, T.57g]. Math.-Nat [T.57h]. Mathematica [New57b]. Mathematical [Ano53i, Ano53k, Ano53l, Ano54g, Ano54h, T.57c]. Mathematics [Goo57b, LWF72, T.57a, T.57e, Gau94, Gau94]. Mathieu [PN58]. Mathison [Tau56]. Matrices [Fro58a, Gre55a, Lot55, MH57, Sch59, Tau50, Wag54, Was52, Wil55, Ano57d]. Matrix [BR53, Doy57, Edm53, FL50, Gre53a, Laa58, Men57a, Opl51, Ril55, SB53, Smi59]. Matter [Ano53a]. Maximum [Lee58, Sad50]. Maximum-Interval [Sad50]. McGraw [Dav57, F.57]. McGraw-Hill [Dav57, F.57]. Mech [New57c]. Mechanical [Luk52]. mekhanikı [Ano57g]. mekhaniky [Ano57h]. Memory [Don52]. Mersenne [Goo55, Leh53, Rie58]. Mesh [Bes54, Gar56]. Method [Bol57, Cam57, CD58, Cur54, DOH54, Dij57, Fle56a, FL50, Gre55a, GS52, Hea57, LaF54, Lot51, Men57b, MH57, Mor56, Mul56, Mul58, Neu59, Pos55, Sha59b, Ski59, Sod53, Spr52, SM58, Vis57, Wag54, WW57, ZS59]. Méthode [T.57b]. Methods [Bes54, Fox51, HH55, Pol59, Wil55]. Might [Sti50]. Million [Tom56]. Milne [Hen57a]. mimeographed [Tom57]. Minimax [GJ59]. Minimum [Bes54, Per52]. Modern [Ano50h]. Modification [Twe54]. Modified [Bla54a, Bla54b, Chi56]. Molecular [Dal54]. Momentum [Mel57]. Monte [DR56, Edm53, FL50, Opl51]. Moscow [Ano57f, Ano57h, Ano57g, Rho57, T.57c]. MTAC [Ano57d, Ano57h, Ano57g, Dia57, T.57f]. multilithed [Hol57b]. Multiple [DR56, HW57, HS58]. Multiplication [Wal58].

N [Hol57b]. N.F [T.57h]. Nachr [New57c]. Nat [T.57h]. naturw [Ewa57a]. Nauk [Ano57f, Ano57h, Ano57g, Rho57]. Naval [Wol52].

NBS [Goo57b, MDL50, T.57e]. Near

[Luk52, Luk56]. Negative [Por56]. Nehari [Hoc58]. Neighborhood [She58]. Neumann [Ano57e]. neunstellige [Ano57d]. Neville [Twe54]. News [Ano50g, Ano53j]. Newton [SK58]. Nine [Ano57d, HSK59]. No [Dia57, Est57, Goo57b, Mel57, T.57e, Tom57]. Nonlinear [Wor55]. NORC [NJ55]. Normal [Har57b, Mul58, Tom56, Vis57, Sch54]. Notation [BWW54]. Notations [Com55]. Note [CD53, Fie58, Gau59, Gib50, T.51, Mil57, Sel53, Sod53, Sti50, Was52, WW57, Boo55, Cle55, Tau50]. notebook [Hol57b]. Notes [Ano50h, Ano57e, ABv^{+}55, AJO^{+}55, Bol57, BRCG55, CR57, CO51, Coh58, DBM55, Dij57, FOP^{+}56, Fox51, FM51, Fro58a, Fro58b, GW56, HR58, HAW56, KR59, Lon58, Lon59, MNR56, Rie58, Rom59, Sal57, SL58, SF59, Ski59, SS59, Sus59, Tau50, Wal58, Wee59, Wil57a, WW57, F.57]. Nouvelle [T.57b]. Nouvelles [Hen57b]. Number [Der59]. Numbers [Coh58, Joh56, Rie58, Rob57, Sha59b, Ano57h, Goo55, Sel53]. Numerical [Ano50h, Bar59, BMv^{+}59, Bla52, Bol57, Cam57, CO51, DF53, F. 50, FGT57, Fet55, Fis57, For53, Fox51, FM51, Goo57a, GL56, Goo57b, Gre57a, GCN59, HMS56, HS56, HW57, HS58, Hen59, Hoc58, HZ56, HPZ59, Hus50, Lon57, Lon58, Luk56, PU59, Pei57, RW59, Sal55, SK58, She55, SZF57, Ste53, SM58, Str57, Tau50, Tom57, Wee59, Wil57a, Wil58, vNG53, Hen57a]. Numérique [Gre57b]. numerischen [New57c].

O [New57c]. O. [Ewa57a, Ewa57b, New57c]. Oberhettinger [F.57]. obituary [Ano57e]. Oblasti [Ano57f]. Obrabotka [Rho57]. Occurring [Dal54]. Off [Ost52]. Office [Goo57b, T.57e, Wol52]. One [Hit57, Est57, Wil57b]. Ontario [Mel57]. Open [Wil57a, Wil58]. Operating [Ele50]. Operations [BW53]. Operators [Bla52, Wil57b]. Optimum [Tha57, Wag57]. Orbits [SZF57]. Order [HS55, HSW56, HSK59, Hit57, Wil57a, Wil58]. Orders [Dia57]. Ordinary [Cam57, Hen59, Ste53, SM58]. Ordnung [New57c]. Organization [Hum54]. Orthogonal [Sal55]. osculatory [T.57g]. ot [Ano57h]. Other [AL51, Gep52, MW51, Max54, Mon50, SR53]. Oxford [T.57a].

p [Ano57d, Ano57h, Ano57g, Coh57, Dia57, Dav57, Ewa57a, Ewa57b, F.57, Goo57b, Gre57b, Hen57a, Hen57b, Hol57b, Mel57, New57a, New57b, New57c, Rho57, T.57c, T57d, T.57f, T.57a, T.57b, T.57e, T.57g]. P. [Sch54, T.57h]. pages [Tom57]. Pair [Smi57]. Papers [ABv^{+}55, AJO^{+}55, BRCG55, CR57, DBM55, FOP^{+}56, GW56, HR58, HAW56, MNR56, SL58, Bol57, Coh58, Dij57, Fro58a, Fro58b, KR59, Lon58, Lon59, Rie58, Rom59, Sal57, SF59, Ski59, SS59, Sus59, Wal58, Wee59, Wil57a, WW57]. Parabolic [DF53]. Parallelströmung [New57c]. Parameter [New57c, T.57h]. Parenthesis [BWW54]. Parenthesis-Free [BWW54]. Paris [Gre57b]. Part [Est57, F.57]. Partial [Fra50, Ho55, Ive53, Sal56]. Partitioning [Wag54]. Partitions [Com55]. Pascal [Cou50]. perekhoda [Ano57h]. Pertaining [BLS57, Hol57b]. Phys [T.57g]. Physik [Ewa57b]. Place [Ano57d, Ano57h]. Places [Rei50a]. Plane [HSW56]. Planes [HSK59]. Planning [F. 50]. Point [GL56, Hen59, Mit59]. Points [Str57]. Poisson [Fet55, Vis57]. polar [Ano57h]. poliarnym [Ano57h]. Polynomial [Atk55, Cle54, Gro54, Hit57, MW59, SR53, Twe54]. Polynomial-Like [Gro54]. Polynomials [CFH58, Dia57, Der59, Men57a, Sal55, Boo55]. population [Sch54]. Positive [Doy57, Ril55]. Positive-Definite [Doy57]. Possibly [Ril55]. Potential [Hoc58]. pound. [Est57]. Power [Fle56b, Men57b]. Powers [Fie58]. Predictors [SS59]. Press [T.57a]. Price [Dav57, Est57, F.57, Goo57b, Gre57b, Hen57a, Hen57b, Mel57, Rho57, T.57c, T.57a, T.57b, T.57e]. Primality [Dij57]. Prime [Smi57]. Primes [KR59, Lee58, Leh53, Sha59a]. Principal [Lon58]. Printing [Goo57b, T.57e]. Probability [Coo57, Sal59]. Problem [Bes54, BMv^{+}59, D. 53, GRS53, HMW55, MDL50, Ril54, Smi57]. Problems [Dal54, GL56, Hen59, Mit59, PU59, War57, Goo57b]. Procedure [Ano53b, Ano54a, Fle56b]. Procedures [Ril54, Twe54]. Proceedings [Est57]. Process [Der59]. Processes [Dav57, Rho57].

Processing [Coh50]. Product [DOH54].

Production [Coo57]. Prog [T57d]. Programmed [MS51, Wag54]. Programming [BD59, Max54, Per52, Wag57]. Programs [Cah54]. Projective [HSW56, HSK59]. Propagation [Tom57, ZS59]. Proper [Men57a, Men57b]. Provision [Ste51]. Pseudo [Joh56]. Publications [Hen57a, T.57b]. Publishing [Hen57b]. Punched [BY51, BY53, Gar54, GS52, Opl51]. Punched-Card [GS52]. Purpose [D. 53, Wal53].

Quadratic [Coh58, Doy57, Sha59a]. Quadrature [DR54, HZ56, HPZ59, Kun56, Luk52, Rei50b, RB59, Sal57, Tha57]. quartic [Coh57].

R [T.57c]. Radio [Tom57]. Radix [Sal51c, Sod53]. Random [Gre55b, Joh56, Mul58, Tom56, Was52, Gru50, GM51, Sch54, Dav57]. Rapid [Lon57]. Rates [Fra50]. Rational [Coo57, Fle56a]. Raymond [Ben50]. REAC [Gep52]. Real [Atk55, GRS53, Men57a, Sou57, T.57f, Coh57]. Rechnungen [New57c]. Recurrence [CO51, GT59, Tod50]. Recurrent [Edm53]. regelmassig [New57c]. Region [Ano57f]. Reihe [Ewa57a]. Related [Hit57]. Relations [CO51, Mor58, Tod50]. Relaxation [Fox51, Gar56, Ost53, Vis57]. Relay [KN51, Wol52]. Remainders [Fro58b]. Remarks [Str57]. Repeated [Kun56]. Report [Cou50, Dia57, Mel57, Tom57]. reprinted [Mel57]. Research [Wol52]. Residues [Sha59a]. Résolution [Gre57b]. Response [Fle56b]. Rev [Ano57g]. Review [Ano57d, Dia57, Hen59, Sch54, T.57c, Tom56]. Reviews [Ano55d]. Richard [Dav57]. Riemann [New57b]. Rigid [GRS53]. River [Mel57]. RMT [Ano57h]. Root [Atk55]. Roots [Atk55, CFH58, Fie58, Gre55a, Wil55, New57b]. Rotation [Hea57]. Rounding [Ost52]. Routines [Hum54]. rubles [T.57c]. Rule [Mon50]. Runge [Lot51].

S [Goo57b, Mel57, T.57e]. S. [Rho57]. Salzer [T.57g]. samples [Sch54]. Sampling [VR51]. Sasaki [Tom57]. Scale [Mul58]. Schrödinger [Ski59]. Schuler [Ano57d]. Scientific [Gar54]. Sea [Tom57]. SEAC [Ele50, Ste51]. Search [GJ59, PN58]. Sechsstellige [T.57h]. Second [Sus59]. See [Ano57h, Ano57g]. Segments [Smi59]. Self [Laa59]. Self-Adjoint [Laa59]. Sequence [F. 50]. Sequences [Rom59]. Sequential [GJ59]. Serial [Don52]. Serial-Memory [Don52]. Series [Cah54, Dav57, Fle56b, Goo57b, Luk59, Mor58, Sal56, T.57e, Cle55]. Set [Cam57, Lot55, Sny55, Ano57d]. Several [Laa59]. Sexagesimal [T.57e]. Shelf [For53]. Shell [Pei57]. Ship [Pol59]. Short [ABv^{+}55, AJO^{+}55, Bol57, BRCG55, CR57, Coh58, DBM55, Dij57, FOP^{+}56, Fro58a, Fro58b, GW56, HR58, HAW56, KR59, Lon58, Lon59, MNR56, Rie58, Rom59, Sal57, SL58, SF59, Ski59, SS59, Sus59, Wal58, Wee59, Wil57a, WW57]. should [T57d]. Sieve [D. 53, Sha59b]. Signals [Lub52]. Simple [Kor52]. Simplex [DOH54]. Simplexes [HMS56, HS56]. Simplified [MH58]. Simulation [GRS53]. Simultaneous [BR53, Laa59, MW51, Sny55]. Single [Ost53, Ost58]. Singularity [Luk52, Luk56]. Size [Gar56]. Skip [Rom59]. Skip-Term [Rom59]. Slide [Mon50]. Small [Gar56, Coh57]. Smoothest [Hol57a]. Snyder [Hol57b]. Soc [Coh57, New57a]. Solution [BR53, BMv^{+}59, Bla52, CO51, F. 50, FGT57, Goo57a, Hen59, Laa58, MW51, Men57b, Mil57, She55, Tod50, Vis57, War57]. Solutions [BLS57, Ber55, HH55, Hol57b, Mit59, Sny55, Ste53, Wor55, de 53]. Solved [Bes54]. Solver [Kor52]. Solving [Atk55, CD58, Cur54, GS52, Mul56, Ril55, Ski59]. Some [Coh58, DR56, Fro58b, NJ55, PU59, RB59, Rob57]. Southard [T.57f]. Special [MH57, Rei50b, Wal53]. Specialized [Sto53]. Specified [Vis57]. Speed [BR53, SA57, Tei52, VR51]. Speeding [Ost58]. Sphere [Coh56]. Spherical [Pei57]. Spherically [Coo57]. Spravochnik [T.57c]. Square [SB53]. SSSR [Ano57f, Ano57h, Ano57g, Rho57]. Stability [Coh56, DF53]. Statistical [MRvN50]. Steiner [HS55]. Steinsalzgitter [Ewa57b]. Steinsalzgitters [Ewa57a]. Step [BM52a, BM52b, Fox51, Her51, Ost53, Ost58, SZF57]. Step-by-Step [Fox51, Her51]. Stirling [SK58]. Stokes [Gre57a]. Storage [Coh50, WR50]. Stream [Gre57a].

Structure [Dal54]. Study

[BMv^{+}59, Est53, vNG53]. Sturm [FGT57]. Sub [Hum54]. Sub-Routines [Hum54]. Subject [Cam57]. Subroutines [MDL50]. Subtraction [Mon50]. Summation [Fet55, Fie58, Rom59, Sal56, Cle55]. Summen [New57c]. Sums [BM52a, BM52b, Ive53, Leh56]. Supplement [Est57]. Swedish [KN51]. Symmetric [BMSW54, Coo57, Doy57, Gre53a, Ril55, Wag54]. Symposium [Gau94]. System [F. 50, Laa59, Pos55]. Systèmes [Gre57b]. Systems [Coh50, Cur54, Dav57, FM52, Gep52, HS55, Ril55, de 53].

T. [T.57f, Tom57]. Tabellen [Ano57d]. Table [Ano57f, Ano57h, Ano57g, BLS57, Dia57, FM51, Lon59, Sal51c, Sou57, T.57f, Hol57b]. Table-Making [FM51]. Tables [Ano53i, Ano53k, Ano53l, Ano54g, Ano54h, Ano55d, Ano57d, Ano57f, Ano57g, Ber55, BY51, BY53, Har57a, Hen57b, Lin57, Lon57, Mel57, Ost52, RW59, Sad50, Sal59, SK58, Sch54, T.57c, Ano57h, Tom57, Ano57d, T.57e, Wil57b]. Tablitsam [T.57c]. Tablitsy [Ano57h, Ano57g, Ano57f, Ano57g]. Tabulation [AS56]. Tafel [T.57h]. Tagung [New57c]. Taylor [Fle56a]. Technical [Ano51f, ABv^{+}55, AJO^{+}55, Boo54, BRCG55, CR57, Coh50, Cou50, DBM55, Don52, Est53, FOP^{+}56, Gar54, GW56, Gre53b, HR58, HAW56, Hum54, Hus50, KN51, MDL50, MNR56, Pea52, SL58, Spr52, Ste51, Sto53, Wal53, Bol57, Coh58, Dij57, Fro58a, Fro58b, KR59, Lon58, Lon59, Rie58, Rom59, Sal57, SF59, Ski59, SS59, Sus59, Wal58, Wee59, Wil57a, WW57]. Technique [Atk55, GRS53, Max54]. Techniques [Est57, GT59]. tekhniki [Ano57h, Ano57g]. Ten [Ano57h]. Term [Rom59]. Terms [Ho55]. Test [Gre55b, Lot55, Ano57d, GM51]. Testing [Joh56]. Tests [Gru50]. their [Sal51c]. Theorem [Sal59]. Theorems [Smi59]. Theory [Ber55, Hoc58, Ost53, Ost58, Wor55, Tom57]. Throwback [Sou56]. Time [GRS53, SZF57, Wal58]. Time-Step [SZF57]. Tochnolı [Ano57g]. Tohnol [Ano57h]. Tokyo [Tom57]. Tompkins [T57d]. Toronto [F.57]. TR [Dia57]. TR-326 [Dia57]. Transcendental [F.57]. Transform [HZ56, HPZ59]. Transformation [Dia57, FM52, Wyn56, Ano57h]. Transforms [Sal55]. Transient [Fle56b, Pol59]. transparency [Wil57b]. Trapezoidal [HH55]. Treatment [DF53, MRvN50]. Treatments [Fra50]. Triangle [Bar59]. Triangular [Wil57b]. Tricomi [F.57]. Trigonometric [LaF54, Sal51c, SR53, Boo55]. Triple [HS55]. Truth [Abb51]. Turing [Tau56]. Two [BMv^{+}59, GL56, Hen59, Leh53, Mit59, Smi59, Hol57b, Wil57b]. Two-Dimensional [BMv^{+}59]. Two-Point [GL56, Hen59]. Type [Cam57, Joh56, Lin57]. typewritten [Wil57b].

U [Dia57, Goo57b, T.57e]. UMT [Hol57b, Wil57b]. Unified [Der59]. Uniqueness [HSW56]. Universitaires [T.57b]. Universitat [Ewa57a]. University [Ano57d, T.57a]. Unpublished [Ano53l]. unterer [New57c]. USAF [Sto53]. Use [BM52a, BM52b, Fox53, Sal59, Tei52, Wil55, GCN59]. Used [SZF57]. Useful [Max54]. Using [Atk55, BWW54, Hoc58, Mul56, SK58, Sny55].

V [T.57c, Ano57d, Ano57f, Ano57h, Ano57g, Coh57, Dia57, Est57, Ewa57a, Ewa57b, F.57, Gre57b, New57a, New57b, New57c, T.57h, T57d, T.57g]. Value [GL56, Mit59]. Values [Lin57, Lon58, Men57a, MRvN50]. Vancouver [Gau94]. Variable [Sal51a]. Vectors [Men57a, Men57b, Wil55]. Verteilter [New57c]. vi [T.57a]. viii [T.57b]. Villars [Gre57b]. Volume [Ano53m]. Volumes [LWF72]. vom [Ewa57b]. vychislitel’noı [Ano57h, Ano57g]. Vyclislitel’nye [Rho57].

W [F.57]. W. [Wil57b]. Walks [Was52]. Wash [T.57e]. Wash- [T.57e]. Washington [Dia57, Goo57b]. Waves [Tom57]. Weierstrass [T.57f, Sou57]. Weight [GCN59]. Weighted [Sal57]. Weights [RW59]. welcher [Ewa57b]. Wert [Ewa57a]. Which [CFH58, Sal51b]. William [Hen57a]. Wilson [Wil57b, Fro58b]. Wiss [T.57h]. Wissenschaftliche [Ewa57a]. without [Her51]. Wolfram [Arc50]. Wolfson [Mel57]. Work [PU59].

X [Boo54]. XIII [Ano50a, Ano50b]. xiv [Hen57b]. xlvi [T.57c]. XV [Ano51a]. xvii [F.57]. xxvi [F.57].

York [Dav57, F.57, Hen57a, Hen57b, T.57a].

Z [Ewa57b, New57c, Ano50a, Ano50b, Ano51a, Ano53c, Ano53d, Ano54b]. zäher [New57c]. zahlentheoretischer [New57c]. Zeitschrift [Ewa57a]. Zero [GJ59, Hit57]. Zeros [Der59, Ive53]. zeta [New57b]. zeta-function [New57b]. Zetafunktionen [New57c]. Zwei [New57c].

References

Abbott:1951:CLT

[Abb51] Wilton R. Abbott. Computing logical truth with the California digital computer. Mathematical Tables and Other Aids to Computation, 5(35):120–128, July 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Archibald:1955:TNSa

[ABv^{+}55] R. C. Archibald, A. D. Booth, John von Neumann, Bryant Tuckerman, B. Zondek, and R. L. La Fara. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 9(49):21–25, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Archibald:1955:TNSb

[AJO^{+}55] R. C. Archibald, Luigi Jacchia, A. M. Ostrowski, Preston C. Hammer, and S. Lipton. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 9(50):62–70, April 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Atkinson:1951:NDM

[AL51] C. P. Atkinson and A. S. Levens. A new differentiating machine (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 5(34):99–102, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1950:BZXa

[Ano50a] Anonymous. Bibliography Z — XIII.

Mathematical Tables and Other Aids to Computation, 4(32):234–238, October 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002501.

Anonymous:1950:BZXb

[Ano50b] Anonymous. Bibliography Z–XIII. Mathematical Tables and Other Aids to Computation, 4(32):239–244, October 1950. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002503.

Anonymous:1950:Ca

[Ano50c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 4 (29):59, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1950:Cb

[Ano50d] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 4 (30):125, April 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1950:Cc

[Ano50e] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 4 (31):180, July 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1950:Cd

[Ano50f] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 4(32):246, October 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1950:N

[Ano50g] Anonymous. News. Mathematical Tables and Other Aids to Computation, 4 (32):238–239, October 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

                                              jstor.org/stable/2002502.

Anonymous:1950:NMN

[Ano50h] Anonymous. Notes on modern numerical analysis—I (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 4(29):39, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1951:BZX

[Ano51a] Anonymous. Bibliography Z–XV. Mathematical Tables and Other Aids to Computation, 5(34):102–108, April 1951. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

                                              jstor.org/stable/2002173.

Anonymous:1951:Ca

[Ano51b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 5 (33):55, January 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1951:Cb

[Ano51c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 5 (34):113, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1951:Cc

[Ano51d] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 5 (35):184, July 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1951:Cd

[Ano51e] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 5(36):259, October 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1951:TDA

[Ano51f] Anonymous. Technical developments (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 5(34):86, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1952:Ca

[Ano52a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 6 (37):61, January 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1952:Cb

[Ano52b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 6 (38):126, April 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1952:Cc

[Ano52c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 6(40):257, October 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1953:BM

[Ano53a] Anonymous. Back matter. Mathematical Tables and Other Aids to Computation, 7 (44):??, October 1953. CODEN MTTCAS.

ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002845.

Anonymous:1953:BCP

[Ano53b] Anonymous. Bibliography of coding procedure. Mathematical Tables and Other Aids to Computation, 7(44):256, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002839.

Anonymous:1953:BZa

[Ano53c] Anonymous. Bibliography Z. Mathematical Tables and Other Aids to Computation, 7(44):256–266, October 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002840.

Anonymous:1953:BZb

[Ano53d] Anonymous. Bibliography Z. Mathematical Tables and Other Aids to Computation, 7(44):268–272, October 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002842.

Anonymous:1953:Ca

[Ano53e] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 7 (41):72, January 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1953:Cb

[Ano53f] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 7 (42):134, April 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1953:Cc

[Ano53g] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 7 (43):213, July 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1953:Cd

[Ano53h] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 7(44):275, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1953:MTE

[Ano53i] Anonymous. Mathematical tables — errata. Mathematical Tables and Other Aids to Computation, 7(44):246–247, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002836.

Anonymous:1953:N

[Ano53j] Anonymous. News. Mathematical Tables and Other Aids to Computation, 7 (44):266–268, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002841.

Anonymous:1953:RMT

[Ano53k] Anonymous. Recent mathematical tables. Mathematical Tables and Other Aids to Computation, 7(44):236–246, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002835.

Anonymous:1953:UMT

[Ano53l] Anonymous. Unpublished mathematical tables. Mathematical Tables and Other Aids to Computation, 7(44):247– 249, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002837.

Anonymous:1953:VI

[Ano53m] Anonymous. Volume information. Mathematical Tables and Other Aids to Computation, 7(44):276–289, October 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

                                              jstor.org/stable/2002831.

Anonymous:1954:BCP

[Ano54a] Anonymous. Bibliography of coding procedure. Mathematical Tables and Other Aids to Computation, 8(48):230–231, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

                                              //www.jstor.org/stable/2002098.

Anonymous:1954:BZa

[Ano54b] Anonymous. Bibliography Z. Mathematical Tables and Other Aids to Computation, 8(48):231–233, October 1954. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

                                              jstor.org/stable/2002099.

Anonymous:1954:Cc

[Ano54c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 8 (47):188, July 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1954:Cd

[Ano54d] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 8(48):245, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1954:Ca

[Ano54e] Anonymous. Corrigendum. Mathematical Tables and Other Aids to Computation, 8 (45):51, January 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1954:Cb

[Ano54f] Anonymous. Corrigendum. Mathematical Tables and Other Aids to Computation, 8 (46):124, April 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1954:MTE

[Ano54g] Anonymous. Mathematical tables—errata.

Mathematical Tables and Other Aids to Computation, 8(48):227–228, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002096.

Anonymous:1954:RMT

[Ano54h] Anonymous. Recent mathematical tables. Mathematical Tables and Other Aids to Computation, 8(48):216–226, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002095.

Anonymous:1955:Ca

[Ano55a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 9 (49):45, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1955:Cb

[Ano55b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 9 (51):137, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1955:Cc

[Ano55c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 9(52):229, October 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1955:RDT

[Ano55d] Anonymous. Reviews and descriptions of tables and books. Mathematical Tables and Other Aids to Computation, 9(51): 126–136, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

                                              tronic). URL http://www.jstor.org/
                                              stable/2002069.

Anonymous:1956:Ca

[Ano56a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 10 (53):55, January 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1956:Cb

[Ano56b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 10 (55):182, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1956:Cc

[Ano56c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 10(56):262–263, October 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1957:Ca

[Ano57a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 11 (57):54, January 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1957:Cc

[Ano57b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 11 (59):228, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1957:Cd

[Ano57c] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 11(60):314, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1957:CHU

[Ano57d] Anonymous. Corrigenda: Harvard University, Computation Laboratory, Annals, Tables of the Function arcsin z, MTAC, v. 10, 1956, Review 66, p. 229, line 3; Mark Lotkin, A set of test matrices, MTAC, v. 9, 1955, p. 161; M. Schuler H. Gebelein,

(a) Acht- und neunstellige Tabellen zu den elliptischen Funktionen (Eight and Nine Place Tables of Elliptical Functions), (b) Fünfstellige Tabellen zu den elliptischen Funktionen (Five Place Tables of Elliptical Functions), MTAC, v. 10, 1956, Review 102, p. 247. Mathematical Tables and Other Aids to Computation, 11(58):129– 130, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). See [Lot55].

Anonymous:1957:NJN

[Ano57e] Anonymous. Notes: John von Neumann 1903–1957 [obituary]. Mathematical Tables and Other Aids to Computation, 11 (58):127, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Anonymous:1957:TEAa

[Ano57f] Anonymous. Table errata: A. A. Abramov, Tablitsy ln Γ[z] v Kompleknoĭ Oblasti (Tables of ln Γ[z] in a Complex Region), Akad. Nauk SSSR, Moscow, 1953. Mathematical Tables and Other Aids to Computation, 11 (58):125, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Anonymous:1957:TEAc

[Ano57g] Anonymous. Table errata: Akademiia Nauk SSSR. [Institut Tochnolĭ mekhanikĭ i vychislitel’noĭ tekhniki]. Matematicheskie Tablitsy. Tablitsy Integralov Frenelia (Tables of Fresnel Integrals), Moscow, 1953. [See Rev. 40, MTAC, v. 9, 1955, p. 75– 76]. Mathematical Tables and Other Aids to Computation, 11(58):126, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

                                              jstor.org/stable/2002157.

Anonymous:1957:TEAb

[Ano57h] Anonymous. Table errata: Akademiia Nauk SSSR. Institut Tohnol mekhaniky i vychislitel’noĭ tekhniki. Matematicheskie Tablitsy. Desiatiéna Tablitsy logarifmov kompleksnykh chisel i perekhoda ot dekartovykh koordinat k poliarnym. Tablitsy funktsiĭ [Ten place tables of logarithms of complex numbers and of the transformation from cartesian to polar coordinates. Tables of functions] ln x, arctg, 2 ln(1+x^{2}), (1+x^{2})^{1/2}. Moscow, 1952. [See RMT 1206, MTAC, v. 8, 1954, p. 149]. Mathematical Tables and Other Aids to Computation, 11(58):125–126, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

                                              jstor.org/stable/2002157.

Anonymous:1958:C

[Ano58] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 12 (62):166, April 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1959:Ca

[Ano59a] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 13 (66):143, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1959:Cb

[Ano59b] Anonymous. Corrigenda. Mathematical Tables and Other Aids to Computation, 13(68):335, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Archibald:1950:NIC

[Arc50] R. C. Archibald. New information concerning Isaac Wolfram’s life and calculations. Mathematical Tables and Other Aids to Computation, 4(32):185–200, October 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Ashour:1956:TF

[AS56] A. Ashour and A. Sabri. Tabulation of the function ψ(θ) = ∑_{n=1}^{∞} \frac{sin nθ}{n^{2}} . Mathematical Tables and Other Aids to Computation, 10(54):57–65, April 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Atkinson:1955:PRS

[Atk55] Cyril Atkinson. Polynomial root solving on the electronic differential analyser (A technique for finding the real and complex roots of a polynomial using an electronic differential analyser). Mathematical Tables and Other Aids to Computation, 9(52):139–143, October 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bartholomew:1959:NIT

[Bar59] Gerald E. Bartholomew. Numerical integration over the triangle. Mathematical Tables and Other Aids to Computation, 13(68):295–298, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bellman:1959:FAD

[BD59] Richard Bellman and Stuart Dreyfus.

Functional approximations and dynamic programming. Mathematical Tables and Other Aids to Computation, 13(68):247– 251, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bennett:1950:RCA

[Ben50] Albert A. Bennett. Raymond Clare Archibald. Mathematical Tables and Other Aids to Computation, 4(29):1–2, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bergman:1955:TDF

[Ber55] Stefan Bergman. Tables for the determination of fundamental solutions of equations in the theory of compressible fluids. Mathematical Tables and Other Aids to Computation, 9(49):8–14, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Best:1954:MPS

[Bes54] G. C. Best. A minimum problem solved by mesh methods. Mathematical Tables and Other Aids to Computation, 8(45):11–13, January 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Brooker:1952:ABA

[BGW52] R. A. Brooker, S. Gill, and D. J. Wheeler. The adventures of a blunder (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 6(38):112–113, April 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Blanch:1952:NSE

[Bla52] Gertrude Blanch. On the numerical solution of equations involving differential operators with constant coefficients. Mathematical Tables and Other Aids to Computation, 6(40):219–223, October 1952. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Blanch:1954:MDDa

[Bla54a] Gertrude Blanch. On modified divided differences I. Mathematical Tables and Other Aids to Computation, 8(45):1–11, January 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Blanch:1954:MDDb

[Bla54b] Gertrude Blanch. On modified divided differences II. Mathematical Tables and Other Aids to Computation, 8(46):67–75, April 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Belford:1957:TPS

[BLS57] G. Belford, L. Jackson Laslett, and J. N.

Snyder. Table pertaining to solutions of a Hill equation. Mathematical Tables and Other Aids to Computation, 11(58): 79–81, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002156.

Brock:1952:UESa

[BM52a] P. Brock and F. J. Murray. The use of exponential sums in step by step integration. Mathematical Tables and Other Aids to Computation, 6(38):63–78, April 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Brock:1952:UESb

[BM52b] P. Brock and F. J. Murray. The use of exponential sums in step by step integration— II. Mathematical Tables and Other Aids to Computation, 6(39):138–150, July 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bivins:1954:CSG

[BMSW54] Robert L. Bivins, N. Metropolis, Paul R. Stein, and Mark B. Wells. Characters of the symmetric groups of degree 15 and 16. Mathematical Tables and Other Aids to Computation, 8(48):212–216, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Blair:1959:SNS

[BMv^{+}59] A. Blair, N. Metropolis, J. von Neumann, A. H. Taub, and M. Tsingou. A study of a numerical solution to a two-dimensional hydrodynamical problem. Mathematical Tables and Other Aids to Computation, 13(67):145–184, July 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Boley:1957:TNS

[Bol57] Bruno A. Boley. Technical Notes and Short Papers: a method for the numerical evaluation of certain infinite integrals. Mathematical Tables and Other Aids to Computation, 11(60):261–264, October 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Booth:1954:TDD

[Boo54] Andrew D. Booth. Technical developments: The development of A.P.E.(X).C. (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 8(46):98–105, April 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Booth:1955:NAP

[Boo55] A. D. Booth. A note on approximating polynomials for trigonometric functions. Mathematical Tables and Other Aids to Computation, 9(49):21–23, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bejarano:1953:SSL

[BR53] Gabriel G. Bejarano and Bruce R. Rosenblatt. A solution of simultaneous linear equations and matrix inversion with high speed computing devices. Mathematical Tables and Other Aids to Computation, 7(42):77–81, April 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Brenner:1955:TNS

[BRCG55] J. L. Brenner, G. W. Reitwiesner, C. W.

Clenshaw, and I. J. Good. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 9(51):117–121, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Brooker:1953:FOE

[BW53] R. A. Brooker and D. J. Wheeler. Floating operations on the EDSAC (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 7(41):37–47, January 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Burks:1954:ALM

[BWW54] Arthur W. Burks, Don W. Warren, and Jesse B. Wright. An analysis of a logical machine using parenthesis-free notation. Mathematical Tables and Other Aids to Computation, 8(46):53–57, April 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Blanch:1951:GTP

[BY51] G. Blanch and E. C. Yowell. A guide to tables on punched cards. Mathematical Tables and Other Aids to Computation, 5(36):185–212, October 1951. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Blanch:1953:AGT

[BY53] Gertrude Blanch and Everett C. Yowell. Addendum to a guide to tables on punched cards. Mathematical Tables and Other Aids to Computation, 7(41):1–6, January 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Cahill:1954:PCH

[Cah54] W. F. Cahill. Programs for computing the hypergeometric series (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 8(45):36–37, January 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Campbell:1957:MIS

[Cam57] Edwin S. Campbell. A method for integrating a set of ordinary differential equations subject to a type of numerical indeterminacy. Mathematical Tables and Other Aids to Computation, 11(60):229–233, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Cahill:1953:NCC

[CD53] W. F. Cahill and Philip Davis. Note: On a computation of the capacity of a cube. Mathematical Tables and Other Aids to Computation, 7(44):272–273, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

                                              //www.jstor.org/stable/2002843.

Conte:1958:ADM

[CD58] S. D. Conte and R. T. Dames. An alternating direction method for solving the biharmonic equation. Mathematical Tables and Other Aids to Computation, 12(63): 198–205, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Campbell:1958:CRP

[CFH58] Edwin S. Campbell, E. M. Fishbach, and J. O. Hirschfelder. Coefficients and roots of the polynomials which define the derivatives of the exponential of (−ϵ/T). Mathematical Tables and Other Aids to Computation, 12(61):1–17, January 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Chisnall:1956:MCE

[Chi56] G. A. Chisnall. A modified Chebyshev– Everett interpolation formula. Mathematical Tables and Other Aids to Computation, 10(54):66–73, April 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clenshaw:1954:PAE

[Cle54] C. W. Clenshaw. Polynomial approximations to elementary functions. Mathematical Tables and Other Aids to Computation, 8(47):143–147, July 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clenshaw:1955:NSC

[Cle55] C. W. Clenshaw. A note on the summation of Chebyshev series. Mathematical Tables and Other Aids to Computation, 9(51): 118–120, 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clenshaw:1951:NNA

[CO51] C. W. Clenshaw and F. W. J. Olver. Notes on numerical analysis—3: Solution of differential equations by recurrence relations (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 5(33):34–39, January 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Cohen:1950:TDM

[Coh50] Arnold A. Cohen. Technical developments: Magnetic drum storage for digital information processing systems (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 4(29): 31–39, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Cohn:1956:SCE

[Coh56] Harvey Cohn. Stability configurations of electrons on a sphere. Mathematical Tables and Other Aids to Computation, 10(55): 117–120, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Cohn:1957:CJG

[Coh57] Harvey Cohn. Comment: J. Godwin, Real quartic fields with small discriminant, London Math. Soc., Jn., v. 31, 1956, p. 478–485. Mathematical Tables and Other Aids to Computation, 11(58):107, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002157.

Cohn:1958:CSB

[Coh58] Harvey Cohn. A computation of some bi-quadratic class numbers (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 12(63):213–217, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Comet:1955:NP

[Com55] Stig Comét. Notations for partitions. Mathematical Tables and Other Aids to Computation, 9(52):143–146, October 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Cook:1957:RFP

[Coo57] J. M. Cook. Rational formulae for the production of a spherically symmetric probability distribution. Mathematical Tables and Other Aids to Computation, 11(58): 81–82, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002156.

Couffignal:1950:TDR

[Cou50] L. Couffignal. Technical developments: Report on the machine of the Institut Blaise Pascal (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 4(32):225–229, October 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Chambers:1957:TNS

[CR57] LL. G. Chambers and Raphael M. Robinson. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 11(57):19–22, January 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Curtiss:1954:GMC

[Cur54] J. H. Curtiss. A generalization of the method of conjugate gradients for solving systems of linear algebraic equations. Mathematical Tables and Other Aids to Computation, 8(48):189–193, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

DHL:1953:SPA

[D. 53] D. H. L. The sieve problem for all-purpose computers. Mathematical Tables and Other Aids to Computation, 7(41):6–14, January 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘D. H. L.’ as author: probably D. H. Lehmer.

Dalgarno:1954:IOP

[Dal54] A. Dalgarno. Integrals occurring in problems of molecular structure. Mathematical Tables and Other Aids to Computation, 8(48):203–212, October 1954. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Davis:1957:CJH

[Dav57] Harold Davis. Comment: J. Halcomb Laning, Jr., and Richard H. Battin, Random Processes in Automatic Control, McGraw-Hill Series in Control Systems Engineering, McGraw-Hill Book Co., Inc., New York, 1956, ix + 434 p., 23 cm. Price $10.00. Mathematical Tables and Other Aids to Computation, 11(58):112, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

Dank:1955:TNS

[DBM55] M. Dank, S. W. Barber, and N. Macon. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 9(52):191–195, October 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

delaGarza:1953:EBA

[de 53] A. de la Garza. Error bounds of approximate solutions to systems of linear algebraic equations. Mathematical Tables and Other Aids to Computation, 7(42):81–84, April 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Derr:1959:UPE

[Der59] John I. Derr. A unified process for the evaluation of the zeros of polynomials over the complex number field. Mathematical Tables and Other Aids to Computation, 13(65): 29–36, January 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

DuFort:1953:SCN

[DF53] E. C. Du Fort and S. P. Frankel. Stability conditions in the numerical treatment of parabolic differential equations. Mathematical Tables and Other Aids to Computation, 7(43):135–152, July 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

DOFL:1957:TEI

[Dia57] Diamond Ordnance Fuze Laboratories. Table errata: Isabelle Arsham, Chebyshev Coefficients for Chebyshev Polynomials of Orders 12 and 24 under the General Linear Transformation (U), Report No. TR-326, Diamond Ordnance Fuze Laboratories, Ordnance Corps, Department of the Army, Washington, D. C., 1956 [Review 72, MTAC, v. 10, 1956, p. 231]. Mathematical Tables and Other Aids to Computation, 11(58):126–127, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

Dijkstra:1957:MIP

[Dij57] E. W. Dijkstra. A method to investigate primality (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 11(59):195–196, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Dantzig:1954:PFI

[DOH54] George B. Dantzig and Wm. Orchard-Hays. The product form for the inverse in the simplex method. Mathematical Tables and Other Aids to Computation, 8(46):64– 67, April 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Donan:1952:TDS

[Don52] John F. Donan. Technical developments:

The serial-memory digital differential analyzer (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 6(38):102–112, April 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Doyle:1957:ISC

[Doy57] Thomas C. Doyle. Inversion of symmetric coefficient matrix of positive-definite quadratic form. Mathematical Tables and Other Aids to Computation, 11(58):55–58, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Davis:1954:EQE

[DR54] P. Davis and P. Rabinowitz. On the estimation of quadrature errors for analytic functions. Mathematical Tables and Other Aids to Computation, 8(48):193–203, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Davis:1956:SMC

[DR56] P. Davis and P. Rabinowitz. Some Monte Carlo experiments in computing multiple integrals. Mathematical Tables and Other Aids to Computation, 10(53):1–8, January 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Edmundson:1953:MCM

[Edm53] H. P. Edmundson. Monte Carlo matrix inversion and recurrent events. Mathematical Tables and Other Aids to Computation, 7(41):18–21, January 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

ELSNBS:1950:OCS

[Ele50] Electronic Laboratory Staff at N. B. S. The operating characteristics of the SEAC (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 4(32):229–230, October 1950. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Estrin:1953:TDE

[Est53] Gerald Estrin. Technical developments:

The electronic computer at the institute for advanced study (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 7(42):108–114, April 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Estrin:1957:CBP

[Est57] Gerald Estrin. Comment: The Proceedings of the Institution of Electrical Engineers, v. 103, Part B, Supplement No. 1– 3, 1956 (Convention on Digital Computer Techniques), British Institution of Electrical Engineers, London. Price one pound. Mathematical Tables and Other Aids to Computation, 11(58):121–124, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002157.

Ewald:1957:CEBa

[Ewa57a] P. P. Ewald. Comment: O. Emersleben, Der Wert der Madelung Konstanten des Steinsalzgitters, Wissenschaftliche Zeitschrift der Universitat Greifswald, v. 3, 1953–54, Math. naturw. Reihe, p. 607– 617. Mathematical Tables and Other Aids to Computation, 11(58):113–114, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002157.

Ewald:1957:CEBb

[Ewa57b] P. P. Ewald. Comment: O. Emersleben, Mit welcher Genauigkeit benötigt man die Madelung-konstante eines Kristalls beispielweise vom Steinsalzgitter? Z. fur Physik. Chemie, v. 204, 1955, p. 43–55. Mathematical Tables and Other Aids to Computation, 11(58):114, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

FJM:1950:PEC

[F. 50] F. J. M. Planning and error considerations for the numerical solution of a system of differential equations on a sequence calculator. Mathematical Tables and Other Aids to Computation, 4(31):133–144, July 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘F. J. M.’ as author; possibly F. J. Murray.

F:1957:CEW

[F.57] A. F. Comment: A. Erd’elyi, W. Magnus, F. Oberhettinger, and F. G. Tricomi, Higher Transcendental Functions. Based, in part, on notes left by H. Bateman. McGraw-Hill, New York, Toronto, London, v. 1, 1953, xxvi + 302 p., 23 cm. Price $6.50; V. 2,1953, xvii + 396 p., 23 cm. Price $7,50; v. 3, 1955, xvii + 292 p., 23 cm. Price $6.50. Mathematical Tables and Other Aids to Computation, 11(58):114–116, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

Fettis:1955:NCC

[Fet55] Henry E. Fettis. Numerical calculation of certain definite integrals by Poisson’s summation formula. Mathematical Tables and Other Aids to Computation, 9(51):85–92, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Farrington:1957:NSS

[FGT57] C. C. Farrington, R. T. Gregory, and A. H. Taub. On the numerical solution of Sturm– Liouville differential equations. Mathematical Tables and Other Aids to Computation, 11(59):131–150, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fielder:1958:NSF

[Fie58] Daniel C. Fielder. A note on summation formulas of powers of roots. Mathematical Tables and Other Aids to Computation, 12(63):194–198, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fishman:1957:NIC

[Fis57] Herbert Fishman. Numerical integration constants. Mathematical Tables and Other Aids to Computation, 11(57):1–9, January 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Forsythe:1950:MIM

[FL50] George E. Forsythe and Richard A.

Leibler. Matrix inversion by a Monte Carlo method. Mathematical Tables and Other Aids to Computation, 4 (31):127–129, July 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL https://www.

ams.org/journals/mcom/1950-04-031/
S0025-5718-1950-0038138-X. Correction

published in volume 5, p. 55 (1951).

Flemming:1956:IMT

[Fle56a] D. P. Flemming. An iterative method for Taylor expansion of rational functions, and applications. Mathematical Tables and Other Aids to Computation, 10(55):120– 130, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Flemming:1956:IPE

[Fle56b] D. P. Flemming. Iterative procedure for evaluating a transient response through its power series. Mathematical Tables and Other Aids to Computation, 10(54):73–81, April 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fox:1951:NNAb

[FM51] L. Fox and J. C. P. Miller. Notes on numerical analysis—5: Table-making for large arguments. the exponential integral (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 5(35):163–166, July 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Forsythe:1952:EGT

[FM52] George E. Forsythe and Theodore S. Motzkin. An extension of Gauss’ transformation for improving the condition of systems of linear equations. Mathematical Tables and Other Aids to Computation, 6(37): 9–17, January 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Forsythe:1956:TNS

[FOP^{+}56] George E. Forsythe, Morio Onoe, Z. Pawlak, R. A. Rubenstein, Marjorie Huse, Stefan Machlup, and Hugh E. Stelson. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 10(53):25–34, January 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Forsythe:1953:NAF

[For53] George E. Forsythe. A numerical analyst’s fifteen-foot shelf. Mathematical Tables and Other Aids to Computation, 7(44):221–228, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fox:1951:NNAa

[Fox51] L. Fox. Notes on numerical analysis— 4: Relaxation and step-by-step methods (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 5(34):92–95, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fox:1953:ULI

[Fox53] L. Fox. The use of large intervals in finitedifference equations. Mathematical Tables and Other Aids to Computation, 7(41):14– 18, January 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Frankel:1950:CRI

[Fra50] Stanley P. Frankel. Convergence rates of iterative treatments of partial differential equations. Mathematical Tables and Other Aids to Computation, 4(30):65–75, April 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Froberg:1958:DHM

[Fro58a] Carl-Erik Froberg. Diagonalization of Hermitian matrices (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 12(63):219– 220, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Froberg:1958:TNS

[Fro58b] Carl-Erik Froberg. Technical Notes and Short Papers: Some computations of Wilson and Fermat remainders. Mathematical Tables and Other Aids to Computation, 12(64):281, October 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Garwick:1954:TDP

[Gar54] Jan V. Garwick. Technical developments: a punched card machine especially equipped for scientific computations (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 8(47): 167–170, July 1954. CODEN MTTCAS.

ISSN 0891-6837 (print), 2326-4853 (electronic).

Garabedian:1956:ERF

[Gar56] P. R. Garabedian. Estimation of the relaxation factor for small mesh size. Mathematical Tables and Other Aids to Computation, 10(56):183–185, October 1956. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gautschi:1959:NBL

[Gau59] Walter Gautschi. Note on bivariate linear interpolation for analytic functions. Mathematical Tables and Other Aids to Computation, 13(66):91–96, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943–1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Greenwood:1959:NIF

[GCN59] R. E. Greenwood, P. D. M. Carnahan, and J. W. Nolley. Numerical integration formulas for use with weight functions x^{2} and x/√\overline{1 − x^{2}}. Mathematical Tables and Other Aids to Computation, 13(65):37–40, January 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gephart:1952:LAS

[Gep52] Landis Gephart. Linear algebraic systems and the REAC (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 6(39):190–203, July 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gibbs:1950:NA

[Gib50] R. C. Gibbs. Note of appreciation. Mathematical Tables and Other Aids to Computation, 4(29):1, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gross:1959:SMS

[GJ59] O. Gross and S. M. Johnson. Sequential minimax search for a zero of a convex function. Mathematical Tables and Other Aids to Computation, 13(65):44–51, January 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Goodman:1956:NIT

[GL56] T. R. Goodman and G. N. Lance. The numerical integration of two-point boundary value problems. Mathematical Tables and Other Aids to Computation, 10(54):82–86, April 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gruenberger:1951:TRD

[GM51] Fred Gruenberger and A. M. Mark. The d^{2} test of random digits. Mathematical Tables and Other Aids to Computation, 5(34): 109–110, 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Good:1955:CCM

[Goo55] I. J. Good. Conjectures concerning the Mersenne numbers. Mathematical Tables and Other Aids to Computation, 9(51):120, September 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Good:1957:NSI

[Goo57a] I. J. Good. On the numerical solution of integral equations. Mathematical Tables and Other Aids to Computation, 11(58): 82–83, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002156.

Goodwin:1957:CNA

[Goo57b] E. T. Goodwin. Comment: NBS Applied Mathematics Series, No. 15, Problems for the Numerical Analysis of the Future, U. S. Gov. Printing Office, Washington, D. C., 1951, iv + 21 p., 26.0 cm. Price $0.20. Mathematical Tables and Other Aids to Computation, 11(58): 116–117, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Gregory:1953:CEE

[Gre53a] Robert T. Gregory. Computing eigenvalues and eigenvectors of a symmetric matrix on the ILLIAC. Mathematical Tables and Other Aids to Computation, 7(44):215–220, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Greig:1953:TDC

[Gre53b] John Greig. Technical developments: The circle computer (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 7(44):249–255, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Greenstadt:1955:MFR

[Gre55a] J. Greenstadt. A method for finding roots of arbitrary matrices. Mathematical Tables and Other Aids to Computation, 9(50):47– 52, April 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Greenwood:1955:CCT

[Gre55b] Robert E. Greenwood. Coupon collector’s test for random digits. Mathematical Tables and Other Aids to Computation, 9(49): 1–5, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002211.

Greenspan:1957:NES

[Gre57a] Donald Greenspan. On the numerical evaluation of the Stokes’ stream function. Mathematical Tables and Other Aids to Computation, 11(59):150–160, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gregory:1957:CCB

[Gre57b] Robert T. Gregory. Comment: L. Couffignal, Résolution Numérique des Systèmes d’Équations Linéaires, Gauthier-Villars, Paris, v. 2, 1956, iii + 180 p., 21 cm. Price $5.86. Mathematical Tables and Other Aids to Computation, 11(58):118– 119, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Gross:1954:PLA

[Gro54] Oliver Gross. Polynomial-like approximation. Mathematical Tables and Other Aids to Computation, 8(46):58–60, April 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gray:1953:TRT

[GRS53] H. J. Gray, Jr., M. Rubinoff, and H. Sohon.

A technique for real time simulation of a rigid body problem. Mathematical Tables and Other Aids to Computation, 7(42):73– 77, April 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gruenberger:1950:TRD

[Gru50] Fred Gruenberger. Tests of random digits. Mathematical Tables and Other Aids to Computation, 4(32):244–245, 1950. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Griffith:1952:PCM

[GS52] B. A. Griffith and K. W. Smillie. On a punched-card method of solving certain integral equations. Mathematical Tables and Other Aids to Computation, 6(39):133–138, July 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Goldstein:1958:BFL

[GT58] M. Goldstein and R. M. Thaler. Bessel functions for large arguments. Mathematical Tables and Other Aids to Computation, 12(61):18–26, January 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Goldstein:1959:RTC

[GT59] M. Goldstein and R. M. Thaler. Recurrence techniques for the calculation of Bessel functions. Mathematical Tables and Other Aids to Computation, 13(66):102– 108, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Goodwin:1956:TNS

[GW56] E. T. Goodwin and P. Wynn. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 10(54):96–100, April 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Harrison:1950:BDC

[Har50] J. O. Harrison, Jr. Binary–decimal conversion on a desk calculator (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 4(32):231–234, October 1950. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002500.

Harris:1957:TEI

[Har57a] Frank E. Harris. Tables of the exponential integral Ei(x). Mathematical Tables and Other Aids to Computation, 11(57):9–16, January 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hart:1957:FAD

[Har57b] Roger G. Hart. A formula for the approximation of definite integrals of the normal distribution function. Mathematical Tables and Other Aids to Computation, 11(60): 265, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Horgan:1956:TNS

[HAW56] R. B. Horgan, E. E. Allen, and D. D. Wall.

Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 10(55):156–168, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Healy:1957:RMC

[Hea57] M. J. R. Healy. A rotation method for computing canonical correlations. Mathematical Tables and Other Aids to Computation, 11(58):83–86, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002156.

Henrici:1957:CAB

[Hen57a] Peter Henrici. Comment: Albert A. Bennett, William E. Milne, and Harry Bateman, Numerical Integration of Differential Equations, Dover Publications, Inc., New York, 1956, 108 p., 20 cm. Price $1.35. Mathematical Tables and Other Aids to Computation, 11(58):117–118, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002157.

Henrici:1957:CDB

[Hen57b] Peter Henrici. Comment: D. Bierens de Haan, Nouvelles Tables d’Intégrales Définies, Hafner Publishing Co., New York, 1957 (edition of 1867-corrected), xiv + 716 p., 24 cm. Price $12.50. Mathematical Tables and Other Aids to Computation, 11 (58):111, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Henrici:1959:BRN

[Hen59] Peter Henrici. Book review: The Numerical Solution of Two-Point Boundary Problems in Ordinary Differential Equations, by L. Fox. Mathematical Tables and Other Aids to Computation, 13(68):322–323, October 1959. CODEN MTTCAS. ISSN

0891-6837. URL https://www.jstor.
org/stable/2002815.

Herrick:1951:SSI

[Her51] Samuel Herrick. Step-by-step integration of ẍ = f(x, y, z, t) without a “corrector”. Mathematical Tables and Other Aids to Computation, 5(34):61–67, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hammer:1955:TMA

[HH55] Preston C. Hammer and Jack W. Hollingsworth. Trapezoidal methods of approximating solutions of differential equations. Mathematical Tables and Other Aids to Computation, 9(51):92–96, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hitchcock:1957:PAB

[Hit57] A. J. M. Hitchcock. Polynomial approximations to Bessel functions of order zero and one and to related functions. Mathematical Tables and Other Aids to Computation, 11(58):86–88, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002156.

Hammer:1956:NISa

[HMS56] P. C. Hammer, O. J. Marlowe, and A. H.

Stroud. Numerical integration over simplexes and cones. Mathematical Tables and Other Aids to Computation, 10(55):130– 137, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Herbst:1955:APC

[HMW55] Eugene H. Herbst, N. Metropolis, and Mark B. Wells. Analysis of problem codes on the MANIAC. Mathematical Tables and Other Aids to Computation, 9(49):14–20, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Ho:1955:DIF

[Ho55] Kuo-Chu Ho. Double interpolation formulae and partial derivatives in terms of finite differences. Mathematical Tables and Other Aids to Computation, 9(50):52–62, April 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hochstrasser:1958:NEP

[Hoc58] U. W. Hochstrasser. Numerical experiments in potential theory using the Nehari estimates. Mathematical Tables and Other Aids to Computation, 12(61):26–33, January 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Holladay:1957:SCA

[Hol57a] John C. Holladay. A smoothest curve approximation. Mathematical Tables and Other Aids to Computation, 11(60):233– 243, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hollander:1957:CGB

[Hol57b] F. H. Hollander. Comment: G. Belford, L. Jackson Laslett, and J. N. Snyder, Table Pertaining to Solutions of a Hill Equation, 1956, 420 p., two 8\frac{1}{2} ×11 multilithed copies in notebook binders deposited in the UMT file. Mathematical Tables and Other Aids to Computation, 11(58):110, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

Hurwitz:1959:NQF

[HPZ59] H. Hurwitz, Jr., R. A. Pfeiffer, and P. F. Zweifel. Numerical quadrature of Fourier transform integrals II. Mathematical Tables and Other Aids to Computation, 13(66): 87–90, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hofsommer:1958:TNS

[HR58] D. J. Hofsommer and Hans Riesel. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 12(61):58–60, January 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hall:1955:DST

[HS55] Marshall Hall, Jr. and J. D. Swift. Determination of Steiner triple systems of order 15. Mathematical Tables and Other Aids to Computation, 9(52):146–152, October 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hammer:1956:NISb

[HS56] Preston C. Hammer and Arthur H.

Stroud. Numerical integration over simplexes. Mathematical Tables and Other Aids to Computation, 10(55):137–139, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hammer:1958:NEM

[HS58] Preston C. Hammer and Arthur H. Stroud.

Numerical evaluation of multiple integrals II. Mathematical Tables and Other Aids to Computation, 12(64):272–280, October 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hall:1959:PPO

[HSK59] Marshall Hall, Jr., J. Dean Swift, and Raymond Killgrove. On projective planes of order nine. Mathematical Tables and Other Aids to Computation, 13(68):233–246, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hall:1956:UPP

[HSW56] Marshall Hall, Jr., J. Dean Swift, and Robert J. Walker. Uniqueness of the projective plane of order eight. Mathematical Tables and Other Aids to Computation, 10(56):186–194, October 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hume:1954:TDI

[Hum54] J. N. P. Hume. Technical developments:

Input and organization of sub-routines for Ferut (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 8(45):30–36, January 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Huskey:1950:TDC

[Hus50] H. D. Huskey. Technical developments: Characteristics of the institute for numerical analysis computer (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 4(30): 103–108, April 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hammer:1957:NEM

[HW57] Preston C. Hammer and A. Wayne Wymore. Numerical evaluation of multiple integrals I. Mathematical Tables and Other Aids to Computation, 11(58):59–67, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Hurwitz:1956:NQF

[HZ56] H. Hurwitz, Jr. and P. F. Zweifel. Numerical quadrature of Fourier transform integrals. Mathematical Tables and Other Aids to Computation, 10(55):140–149, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Isbell:1959:EMG

[Isb59] John R. Isbell. On the enumeration of majority games. Mathematical Tables and Other Aids to Computation, 13(65):21–28, January 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Iverson:1953:ZPS

[Ive53] K. E. Iverson. The zeros of the partial sums of e^{z}. Mathematical Tables and Other Aids to Computation, 7(43):165–168, July 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Johnson:1956:GTP

[Joh56] D. L. Johnson. Generating and testing pseudo random numbers on the IBM Type 701. Mathematical Tables and Other Aids to Computation, 10(53):8–13, January 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Kjellberg:1951:TDB

[KN51] Goran Kjellberg and Gosta Neovius. Technical developments: The BARK, A Swedish general relay computer (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 5(33):29–34, January 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Korn:1952:DAS

[Kor52] Granino A. Korn. The difference analyzer:

a simple differential equation solver. Mathematical Tables and Other Aids to Computation, 6(37):1–8, January 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Killgrove:1959:CCP

[KR59] R. B. Killgrove and K. E. Ralston. On a conjecture concerning the primes (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(66):121–122, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Kunz:1956:HAQ

[Kun56] K. S. Kunz. High accuracy quadrature formulas from divided differences with repeated arguments. Mathematical Tables and Other Aids to Computation, 10(54): 87–90, April 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Laasonen:1958:ISM

[Laa58] Pentti Laasonen. On the iterative solution of the matrix equation AX^{2} −I = 0. Mathematical Tables and Other Aids to Computation, 12(62):109–116, April 1958. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Laasonen:1959:SDS

[Laa59] P. Laasonen. On the simultaneous determination of several eigensolutions of a self-adjoint system of differential equations. Mathematical Tables and Other Aids to Computation, 13(65):13–20, January 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

LaFara:1954:MCI

[LaF54] Robert L. LaFara. A method for calculating inverse trigonometric functions. Mathematical Tables and Other Aids to Computation, 8(47):132–139, July 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Leech:1958:GPH

[Lee58] John Leech. Groups of primes having maximum density. Mathematical Tables and Other Aids to Computation, 12(62):144– 145, April 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lehmer:1953:TNM

[Leh53] D. H. Lehmer. Two new Mersenne primes. Mathematical Tables and Other Aids to Computation, 7(41):72, January 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lehmer:1956:LGS

[Leh56] Emma Lehmer. On the location of Gauss sums. Mathematical Tables and Other Aids to Computation, 10(56):194–202, October 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Ling:1957:TVI

[Lin57] Chih-Bing Ling. Tables of values of 16 integrals of algebraic-hyperbolic type. Mathematical Tables and Other Aids to Computation, 11(59):160–166, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Longman:1957:TRA

[Lon57] I. M. Longman. Tables for the rapid and accurate numerical evaluation of certain infinite integrals involving Bessel functions. Mathematical Tables and Other Aids to Computation, 11(59):166–180, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Longman:1958:NEC

[Lon58] I. M. Longman. On the numerical evaluation of Cauchy principal values of integrals (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 12(63):205–207, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Longman:1959:STT

[Lon59] I. M. Longman. A short table of ∫_{x}^{∞} J_{0}(t)t^{−n}dt and ∫_{x}^{∞} J_{1}(t)t^{−n}dt (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(68):306–311, October 1959. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lotkin:1951:ARK

[Lot51] Max Lotkin. On the accuracy of Runge– Kutta’s method. Mathematical Tables and Other Aids to Computation, 5(35):128–133, July 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lotkin:1955:STM

[Lot55] Mark Lotkin. A set of test matrices. Mathematical Tables and Other Aids to Computation, 9(52):153–161, October 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). See corrigendum [Ano57d].

Lubkin:1952:ASD

[Lub52] S. Lubkin. Asynchronous signals in digital computers (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 6 (40):238–241, October 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Luke:1952:MQN

[Luk52] Yudell L. Luke. Mechanical quadrature near a singularity. Mathematical Tables and Other Aids to Computation, 6 (40):215–219, October 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Luke:1956:EIA

[Luk56] Yudell L. Luke. Evaluation of an integral arising in numerical integration near a logarithmic singularity. Mathematical Tables and Other Aids to Computation, 10(53): 14–21, January 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Luke:1957:C

[Luk57] Yudell L. Luke. On the computation of log Z and arc tan Z. Mathematical Tables and Other Aids to Computation, 11(57): 16–18, January 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Luke:1959:ECH

[Luk59] Yudell L. Luke. Expansion of the confluent hypergeometric function in series of Bessel functions. Mathematical Tables and Other Aids to Computation, 13(68):261–271, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Luke:1972:CIM

[LWF72] Yudell L. Luke, Jet Wimp, and Wyman Fair, editors. Cumulative Index to Mathematics of Computation, Volumes 1–24 (1943–1969). American Mathematical Society, Providence, RI, USA, 1972. xvii + 461 pp.

Maxfield:1954:UTP

[Max54] J. E. Maxfield. A useful technique in programming for analog computers (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 8 (48):233–234, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

MDLStaff:1950:TDI

[MDL50] MDL Staff. Technical developments: The incorporation of subroutines into a complete problem on the NBS eastern automatic computer (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 4(31):164–168, July 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Melkanoff:1957:CJW

[Mel57] M. A. Melkanoff. Comment: J. L. Wolfson and H. S. Gellman, Five Figure Tables for the Conversion of Electron Momentum to Electron Kinetic Energy from 100 to 20000 Gauss-Cm, Atomic Energy of Canada, Ltd., Chalk River, Ontario, 1954 (reprinted July, 1956), A.E.C.L. Report No. 327, i + 43 p., 27 cm. Price $1.00. Mathematical Tables and Other Aids to Computation, 11(58):112–113, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

Mendelsohn:1957:CCP

[Men57a] N. S. Mendelsohn. The computation of complex proper values and vectors of a real matrix with application to polynomials. Mathematical Tables and Other Aids to Computation, 11(58):91–94, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

                                              //www.jstor.org/stable/2002156.

Mendelsohn:1957:IMS

[Men57b] N. S. Mendelsohn. An iterative method for the solution of linear equations based on the power method for proper vectors. Mathematical Tables and Other Aids to Computation, 11(58):88–91, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

                                              //www.jstor.org/stable/2002156.

Meyer:1957:MIL

[MH57] H. I. Meyer and B. J. Hollingsworth. A method of inverting large matrices of special form. Mathematical Tables and Other Aids to Computation, 11(58):94–97, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

                                              //www.jstor.org/stable/2002156.

Miller:1958:SCE

[MH58] James Miller and R. P. Hurst. Simplified calculation of the exponential integral. Mathematical Tables and Other Aids to Computation, 12(63):187–193, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Miller:1950:CD

[Mil50] J. C. P. Miller. Checking by differences— I. Mathematical Tables and Other Aids to Computation, 4(29):3–11, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Miller:1957:NGS

[Mil57] J. C. P. Miller. Note on the general solution of the confluent hypergeometric equation. Mathematical Tables and Other Aids to Computation, 11(58):97–99, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002156.

Mirsepassi:1959:GEC

[Mir59] T. Mirsepassi. Graphical evaluation of a convolution integral. Mathematical Tables and Other Aids to Computation, 13(67): 202–212, July 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Mitchell:1959:ICB

[Mit59] A. R. Mitchell. The influence of critical boundary conditions on finite difference solutions of two point boundary value problems. Mathematical Tables and Other Aids to Computation, 13(68):252–260, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Morton:1956:TNS

[MNR56] K. W. Morton, H. Nagler, and Hans Riesel.

Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 10(56):223–227, October 1956. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Montgomerie:1950:ASL

[Mon50] G. A. Montgomerie. Addition and subtraction on a logarithmic slide rule (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 4(29): 53, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Morton:1951:CDC

[Mor51] Paul L. Morton. The California digital computer. Mathematical Tables and Other Aids to Computation, 5(34):57–61, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Morrison:1956:MCC

[Mor56] D. R. Morrison. A method for computing certain inverse functions. Mathematical Tables and Other Aids to Computation, 10(56):202–208, October 1956. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Moran:1958:ARB

[Mor58] P. A. P. Moran. Approximate relations between series and integrals. Mathematical Tables and Other Aids to Computation, 12(61):34–37, January 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Metropolis:1950:STV

[MRvN50] N. C. Metropolis, G. Reitwiesner, and J. von Neumann. Statistical treatment of values of first 2, 000 decimal digits of e and of π calculated on the ENIAC. Mathematical Tables and Other Aids to Computation, 4(30):109–111, 1950.

MacMillan:1951:FDC

[MS51] Donald B. MacMillan and Richard H. Stark. “floating decimal” calculation on the IBM card programmed electronic calculator (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 5(34):86–92, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Muller:1956:MSA

[Mul56] David E. Muller. A method for solving algebraic equations using an automatic computer. Mathematical Tables and Other Aids to Computation, 10(56):208–215, October 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Muller:1958:IMG

[Mul58] Mervin E. Muller. An inverse method for the generation of random normal deviates on large-scale computers. Mathematical Tables and Other Aids to Computation, 12(63):167–174, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Maxfield:1951:GSS

[MW51] John E. Maxfield and Frederick V. Waugh.

A graphic solution of simultaneous linear equations (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 5(36):246–248, October 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Murnaghan:1959:DCA

[MW59] F. D. Murnaghan and J. W. Wrench, Jr.

The determination of the Chebyshev approximating polynomial for a differentiable function. Mathematical Tables and Other Aids to Computation, 13(67):185–193, July 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Neustadt:1959:MCE

[Neu59] Lucien Neustadt. A method of computing eigenvectors and eigenvalues on an analog computer. Mathematical Tables and Other Aids to Computation, 13(67):194–201, July 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Newman:1957:CDHa

[New57a] Morris Newman. Comment: D. H. Lehmer, On the diophantine equation x^{3} + y^{3} + z^{3} = 1, London Math. Soc., Jn., v. 31, 1956, p. 275–280. Mathematical Tables and Other Aids to Computation, 11(58): 106–107, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Newman:1957:CDHb

[New57b] Morris Newman. Comment: D. H. Lehmer, On the roots of the Riemann zeta-function, Acta Mathematica, v. 95, 1956, p. 291–298. Mathematical Tables and Other Aids to Computation, 11(58): 107–108, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Newman:1957:CEB

[New57c] Morris Newman. Comment: O. Emersleben, Anwendungen zahlentheoretischer Abschatzungen bei numerischen Rechnungen, Z. angew. Math. Mech., v. 33, 1953, p. 1-3. O. Emersleben, Über Summen Epsteinscher Zetafunktionen regelmassig Verteilter unterer’ Parameter, Math. Nachr., v. 13, 1955, p. 59-72. O. Emersleben, Über Zwei Epsteinsche Zetafunktionen 4. und 8. Ordnung, Ber. Math. Tagung, 1953, p. 233-250. O. Emersleben, Über eine doppelperiodische Parallelströmung zäher Flüssigkeiten, Z. angew. Math. Mech., v. 35, 1955, p. 156–160. Mathematical Tables and Other Aids to Computation, 11(58): 109–110, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Nicholson:1955:SCN

[NJ55] S. C. Nicholson and J. Jeenel. Some comments on a NORC computation of π. Mathematical Tables and Other Aids to Computation, 9(52):162–164, October 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002052.

Opler:1951:MCM

[Opl51] Ascher Opler. Monte Carlo matrix calculation with punched card machines. Mathematical Tables and Other Aids to Computation, 5(35):115–120, July 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Ostrowski:1952:RDT

[Ost52] A. M. Ostrowski. On the rounding off of difference tables for linear interpolation. Mathematical Tables and Other Aids to Computation, 6(40):212–214, October 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Ostrowski:1953:URT

[Ost53] A. Ostrowski. On over and under relaxation in the theory of the cyclic single step iteration. Mathematical Tables and Other Aids to Computation, 7(43):152–159, July 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Ostrowski:1958:GSD

[Ost58] Alexander M. Ostrowski. On Gauss’ speeding up device in the theory of single step iteration. Mathematical Tables and Other Aids to Computation, 12(62):116– 132, April 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Pearcey:1952:TDA

[Pea52] T. Pearcey. Technical developments: An automatic computer in Australia (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 6(39):167–172, July 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Peirce:1957:NIS

[Pei57] William H. Peirce. Numerical integration over the spherical shell. Mathematical Tables and Other Aids to Computation, 11(60):244–249, October 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Perry:1952:MAP

[Per52] David P. Perry. Minimum access programming (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 6(39): 172–182, July 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Parker:1958:SAM

[PN58] E. T. Parker and Paul J. Nikolai. A search for analogues of the Mathieu group. Mathematical Tables and Other Aids to Computation, 12(61):38–43, January 1958. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Polachek:1959:CTE

[Pol59] Harry Polachek. Calculation of transient excitation of ship hulls by finite difference methods. Mathematical Tables and Other Aids to Computation, 13(66):109– 116, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Porter:1956:IND

[Por56] R. J. Porter. On irregular negative determinants of exponent 9n. Mathematical Tables and Other Aids to Computation, 10(53): 22–25, January 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Postley:1955:MES

[Pos55] J. A. Postley. A method for the evaluation of a system of Boolean algebraic equations. Mathematical Tables and Other Aids to Computation, 9(49):5–8, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Pasta:1959:HNW

[PU59] John R. Pasta and S. Ulam. Heuristic numerical work in some problems of hydrodynamics. Mathematical Tables and Other Aids to Computation, 13(65):1–12, January 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Riley:1959:GQS

[RB59] J. A. Riley and C. Billings. Gaussian quadrature of some integrals involving Airy functions. Mathematical Tables and Other Aids to Computation, 13(66):97–101, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Reitwiesner:1950:EDM

[Rei50a] George W. Reitwiesner. An ENIAC determination of π and e to more than 2000 decimal places. Mathematical Tables and Other Aids to Computation, 4(29):11–15, January 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002695.

Reiz:1950:SCQ

[Rei50b] A. Reiz. On a special case of a quadrature formula of christoffel. Mathematical Tables and Other Aids to Computation, 4(32):181–185, October 1950. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Rhodes:1957:CLE

[Rho57] Ida Rhodes. Comment: S. A. Lebedev, Elektronnye Vyclislitel’nye Mashiny i Obrabotka Informatksi (Electronic Computing Machines and Information Processes), Akad. Nauk SSSR, Moscow, 1956, 47 p., 20 cm. Price 65 kopeks. Mathematical Tables and Other Aids to Computation, 11(58):124–125, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

Riesel:1958:MNT

[Rie58] Hans Riesel. Mersenne numbers (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 12(63):207–213, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Riley:1954:IPD

[Ril54] James D. Riley. Iteration procedures for the Dirichlet difference problem. Mathematical Tables and Other Aids to Computation, 8(47):125–131, July 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Riley:1955:SSL

[Ril55] James D. Riley. Solving systems of linear equations with a positive definite, symmetric, but possibly ill-conditioned matrix. Mathematical Tables and Other Aids to Computation, 9(51):96–101, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Ritchie:1950:DI

[Rit50] R. H. Ritchie. On a definite integral. Mathematical Tables and Other Aids to Computation, 4(30):75–77, April 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Robinson:1957:SFN

[Rob57] Raphael M. Robinson. Some factorizations of numbers of the form 2^{n} ± 1. Mathematical Tables and Other Aids to Computation, 11(60):265–268, October 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Roman:1959:STS

[Rom59] Irwin Roman. Skip-term summation of sequences (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(66):117–121, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Rabinowitz:1959:TAW

[RW59] Philip Rabinowitz and George Weiss. Tables of abscissas and weights for numerical evaluation of integrals of the form ∫_{0}^{∞} e^{−x}x^{n}f(x)dx. Mathematical Tables and Other Aids to Computation, 13 (68):285–294, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stegun:1957:GBF

[SA57] Irene A. Stegun and Milton Abramowitz.

Generation of Bessel functions on high speed computers. Mathematical Tables and Other Aids to Computation, 11(60):255– 257, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Sadler:1950:MIT

[Sad50] D. H. Sadler. Maximum-interval tables. Mathematical Tables and Other Aids to Computation, 4(31):129–132, July 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1951:FCE

[Sal51a] H. E. Salzer. Formulas for calculating the error function of a complex variable. Mathematical Tables and Other Aids to Computation, 5(34):67–70, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1951:FFA

[Sal51b] H. E. Salzer. Formulas for finding the argument for which a function has a given derivative. Mathematical Tables and Other Aids to Computation, 5(36):213–215, October 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1951:RTT

[Sal51c] H. E. Salzer. Radix table for trigonometric functions and their inverses to high accuracy. Mathematical Tables and Other Aids to Computation, 5(33):9–11, January 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1955:OPA

[Sal55] Herbert E. Salzer. Orthogonal polynomials arising in the numerical evaluation of inverse Laplace transforms. Mathematical Tables and Other Aids to Computation, 9(52):164–177, October 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1956:FPS

[Sal56] Herbert E. Salzer. Formulas for the partial summation of series. Mathematical Tables and Other Aids to Computation, 10(55): 149–156, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1957:EWQ

[Sal57] Herbert E. Salzer. Equally-weighted quadrature formulas for inversion integrals (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 11(59):197–200, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salton:1959:UCL

[Sal59] Gerard Salton. The use of the central limit theorem for interpolating in tables of probability distribution functions. Mathematical Tables and Other Aids to Computation, 13(67):213–216, July 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Saibel:1953:FCE

[SB53] Edward Saibel and W. J. Berger. On finding the characteristic equation of a square matrix. Mathematical Tables and Other Aids to Computation, 7(44):228–236, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Schwartz:1954:RPC

[Sch54] Lorraine Schwartz. Review of P. C. Mahalanobis, Tables of random samples from a normal population. Mathematical Tables and Other Aids to Computation, 8 (48):228–230, October 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Schechter:1959:ICM

[Sch59] Samuel Schechter. On the inversion of certain matrices. Mathematical Tables and Other Aids to Computation, 13(66):73–77, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Selfridge:1953:NFF

[Sel53] J. L. Selfridge. Note: Factors of Fermat numbers. Mathematical Tables and Other Aids to Computation, 7(44):274– 275, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002843.

Sherry:1959:CGF

[SF59] M. E. Sherry and S. Fulda. Calculation of gamma functions to high accuracy (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(68):314–315, October 1959. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Shanks:1954:LA

[Sha54] Daniel Shanks. A logarithm algorithm.

Mathematical Tables and Other Aids to Computation, 8(46):60–64, April 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Shanks:1959:QRD

[Sha59a] Daniel Shanks. Quadratic residues and the distribution of primes. Mathematical Tables and Other Aids to Computation, 13(68):272–284, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Shanks:1959:SMF

[Sha59b] Daniel Shanks. A sieve method for factoring numbers of the form n^{2} + 1. Mathematical Tables and Other Aids to Computation, 13(66):78–86, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Sheldon:1955:NSE

[She55] John W. Sheldon. On the numerical solution of elliptic difference equations. Mathematical Tables and Other Aids to Computation, 9(51):101–112, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Sheldon:1958:AAL

[She58] J. W. Sheldon. Algebraic approximations for Laplace’s equation in the neighborhood of interfaces. Mathematical Tables and Other Aids to Computation, 12(63):174– 186, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1958:ELT

[SK58] Herbert E. Salzer and Genevieve M. Kimbro. Extension of Lindow’s tables for numerical differentiation using Newton– Stirling and Newton–Bessel differences. Mathematical Tables and Other Aids to Computation, 12(62):133–140, April 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Skillman:1959:EMS

[Ski59] Sherwood Skillman. Efficient method for solving atomic Schrödinger’s equation (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(68):299–302, October 1959. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1958:TNS

[SL58] Herbert E. Salzer and Norman Levine.

Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 12(62):141–144, April 1958. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stoller:1958:MNI

[SM58] L. Stoller and D. Morrison. A method for the numerical integration of ordinary differential equations. Mathematical Tables and Other Aids to Computation, 12 (64):269–272, October 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Smith:1957:GPP

[Smi57] Herschel F. Smith. On a generalization of the prime pair problem. Mathematical Tables and Other Aids to Computation, 11(60):249–254, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Smith:1959:TTI

[Smi59] Richard B. Smith. Two theorems on inverses of finite segments of the generalized Hilbert matrix. Mathematical Tables and Other Aids to Computation, 13(65):41–43, January 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Snyder:1955:ISS

[Sny55] James N. Snyder. On the improvement of the solutions to a set of simultaneous linear equations using the ILLIAC. Mathematical Tables and Other Aids to Computation, 9(52):177–184, October 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Soden:1953:NMR

[Sod53] Walter Soden. Note: a method of radix conversion. Mathematical Tables and Other Aids to Computation, 7(44):273– 274, October 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002843.

Southard:1956:EFB

[Sou56] Thomas H. Southard. Everett’s formula for bivariate interpolation and throwback of fourth differences. Mathematical Tables and Other Aids to Computation, 10 (56):216–223, October 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Southard:1957:ATW

[Sou57] Thomas H. Southard. Approximation and table of the Weierstrass ℘ function in the equianharmonic case for real argument. Mathematical Tables and Other Aids to Computation, 11(58):99–100, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002156.

Sprague:1952:TDF

[Spr52] R. E. Sprague. Technical developments: Fundamental concepts of the digital differential analyzer method of computation (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 6(37):41–49, January 1952. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Scherberg:1953:ACP

[SR53] Max G. Scherberg and John F. Riordan.

Analogue calculation of polynomial and trigonometric expansions (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 7(41):61–65, January 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stoller:1959:CCC

[SS59] D. S. Stoller and L. C. Stoller. Calculating the coefficients of certain linear predictors (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(66):122– 124, April 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stein:1951:TDP

[Ste51] P. Stein. Technical developments: Provision for expansion in the SEAC (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 5(36):232–240, October 1951. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Sterne:1953:ANS

[Ste53] Theodore E. Sterne. The accuracy of numerical solutions of ordinary differential equations. Mathematical Tables and Other Aids to Computation, 7(43):159–164, July 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stibitz:1950:NMC

[Sti50] G. R. Stibitz. A note on “is” and “might be” in computers (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 4(31): 168–169, July 1950. CODEN MTTCAS.

ISSN 0891-6837 (print), 2326-4853 (electronic).

Stone:1953:TDU

[Sto53] J. J. Stone. Technical developments: The USAF–Fairchild Specialized Computer (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 7(41):34–37, January 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stroud:1957:RDP

[Str57] A. H. Stroud. Remarks on the disposition of points in numerical integration formulas. Mathematical Tables and Other Aids to Computation, 11(60):257–261, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Suschowk:1959:EFA

[Sus59] Dietrich Suschowk. Explicit formulae for 25 of the associated Legendre functions of the second kind (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(68):303–305, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Sheldon:1957:TSU

[SZF57] J. W. Sheldon, B. Zondek, and M. Friedman. On the time-step to be used for the computation of orbits by numerical integration. Mathematical Tables and Other Aids to Computation, 11(59):181–189, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

JT:1951:ENA

[T.51] J. T. Editorial note (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 5(35):167, July 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). Original journal has only ‘J. T.’ as author.

T:1957:CJC

[T.57a] A. H. T. Comment: J. Crank, The Mathematics of Diffusion, Oxford University Press, New York, 1956, vi + 347 p., 23.5 cm. Price $8.00. Mathematical Tables and Other Aids to Computation, 11(58): 119–120, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

T:1957:CMV

[T.57b] A. H. T. Comment: Marcel van Laethem, Une Méthode Nouvelle et Générale de Calcul des Intégrales Géneralisiées, Publications Universitaires de Louvain, Louvain (Belgique), 1956, viii + 180 p., 23 cm. Price 250 Fr.b. ($5.00). Mathematical Tables and Other Aids to Computation, 11(58): 120–121, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

T:1957:BRV

[T.57c] C. B. T. Book review: A. V. Lebedev and R. M. Fedorova, Spravochnik po Matematicheskim Tablitsam (Guide to Mathematical Tables), Moscow, 1956, xlvi + 549 p., 26 cm. Price 29.20 rubles. Mathematical Tables and Other Aids to Computation, 11(58):104–106, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

T:1957:CFH

[T57d] C. B. T. Comment: F. H. Hollander and C. B. Tompkins, What you should know about digital computers, Chem. Eng. Prog., v. 52, 1956, p. 451–454. Mathematical Tables and Other Aids to Computation, 11 (58):110, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

                                              tronic). URL http://www.jstor.org/
                                              stable/2002157.

T:1957:CNA

[T.57e] C. B. T. Comment: NBS Applied Mathematics Series, No. 35, Tables of Lagrangian Coefficients for Sexagesimal Interpolation, U. S. Govt. Printing Office, Wash- ington, D. C., 1954, ix + 157 p., 26 cm. Price $2.00. Mathematical Tables and Other Aids to Computation, 11(58):108, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

                                              jstor.org/stable/2002157.

T:1957:CHS

[T.57f] C. B. T. Comment: T. H. Southard, Approximation and table of the Weierstrass ℘ function in the equianharmonic case for real argument. [MTAC, this issue, p. 99–100]. Mathematical Tables and Other Aids to Computation, 11(58):110, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

                                              //www.jstor.org/stable/2002157.

T:1957:CSB

[T.57g] J. T. Comment: E. Salzer, Coefficients for complex osculatory interpolation over a Cartesian grid, J. Math. Phys., v. 35, 1956, p. 152–163. Mathematical Tables and Other Aids to Computation, 11(58):109, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

                                              //www.jstor.org/stable/2002157.

T:1957:CEG

[T.57h] J. T. Comment: Ernst Glowatzki, Sechsstellige Tafel der Cauer-Parameter, Abh. Bayer. Akad. Wiss. Math.-Nat. Kl. (N.F.), v. 67, 1955, 37 p. Mathematical Tables and Other Aids to Computation, 11 (58):111, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.jstor.org/
stable/2002157.

Taussky:1950:NNA

[Tau50] Olga Taussky. Notes on numerical analysis. II. Note on the condition of matrices. Mathematical Tables and Other Aids to Computation, 4(30):111–112, 1950. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Taussky:1956:AMT

[Tau56] Olga Taussky. Alan Mathison Turing 1912–1954. Mathematical Tables and Other Aids to Computation, 10(55):180– 181, 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Teichroew:1952:UCF

[Tei52] D. Teichroew. Use of continued fractions in high speed computing. Mathematical Tables and Other Aids to Computation, 6(39): 127–133, July 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Thacher:1957:OQF

[Tha57] Henry C. Thacher, Jr. Optimum quadrature formulas in s dimensions. Mathematical Tables and Other Aids to Computation, 11(59):189–194, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Todd:1950:SDE

[Tod50] John Todd. Solution of differential equations by recurrence relations (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 4(29):39–44, January 1950. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Tompkins:1956:RMR

[Tom56] C. B. Tompkins. Review of A Million Random Digits With 100,000 Normal Deviates. Mathematical Tables and Other Aids to Computation, 10(53):39–43, January 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Tomizuka:1957:CSB

[Tom57] C. T. Tomizuka. Comment: T. Sasaki, The Effect of Earth and Sea on the Propagation of Radio Waves, Numerical Computation Bureau, Report No. 6, Tokyo, 1952. 14 mimeographed pages of theory, 25 cm., and 6 tables of functions, 25 × 34 cm. Mathematical Tables and Other Aids to Computation, 11(58):111–112, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.

jstor.org/stable/2002157.

Tweedie:1954:MAN

[Twe54] M. C. K. Tweedie. A modification of the Aitken–Neville linear iterative procedures for polynomial interpolation. Mathematical Tables and Other Aids to Computation, 8(45):13–16, January 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Viswanathan:1957:SPE

[Vis57] R. V. Viswanathan. Solution of Poisson’s equation by relaxation method— normal gradient specified on curved boundaries. Mathematical Tables and Other Aids to Computation, 11(58):67–78, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

vonNeumann:1953:NSC

[vNG53] J. von Neumann and H. H. Goldstine. A numerical study of a conjecture of Kummer. Mathematical Tables and Other Aids to Computation, 7:133–134, 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

vonNeumann:1955:CFE

[vNT55] John von Neumann and Bryant Tuckerman. Continued fraction expansion of 2^{1/3}. Mathematical Tables and Other Aids to Computation, 9(49):23–25, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Votaw:1951:HSS

[VR51] D. F. Votaw, Jr. and J. A. Rafferty.

High speed sampling. Mathematical Tables and Other Aids to Computation, 5(33):1–8, January 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wagner:1954:PMI

[Wag54] Harvey M. Wagner. A partitioning method of inverting symmetric definite matrices on a card-programmed calculator. Mathematical Tables and Other Aids to Computation, 8(47):139–143, July 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wagner:1957:LAO

[Wag57] Harvey M. Wagner. A lemma for automatic optimum programming on the IBM 650. Mathematical Tables and Other Aids to Computation, 11(58):101–104, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002156.

Walker:1953:TDS

[Wal53] J. P. Walker, Jr. Technical developments:

a special purpose digital computer (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 7(43):190–195, July 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wall:1958:MTI

[Wal58] D. D. Wall. Multiplication time on the IBM 709 (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 12(63):217–218, July 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Warner:1957:SJP

[War57] F. J. Warner. On the solution of “jury” problems with many degrees of freedom. Mathematical Tables and Other Aids to Computation, 11(60):268–271, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wasow:1952:NIM

[Was52] W. R. Wasow. A note on the inversion of matrices by random walks. Mathematical Tables and Other Aids to Computation, 6(38):78–81, April 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wasow:1958:AID

[Was58] Wolfgang Wasow. On the accuracy of implicit difference approximations to the equation of heat flow. Mathematical Tables and Other Aids to Computation, 12(61): 43–55, January 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Weeg:1959:NIT

[Wee59] Gerard P. Weeg. Numerical integration of

∫_{0}^{∞} e^{−x}J_{0}(\frac{ηx}{ξ})J_{1}(\frac{x}{ξ})x^{−n}dx (in Techni-

cal Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(68):312–313, October 1959. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wilkinson:1955:UIM

[Wil55] J. H. Wilkinson. The use of iterative methods for finding the latent roots and vectors of matrices. Mathematical Tables and Other Aids to Computation, 9(52):184–191, October 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wilf:1957:OFN

[Wil57a] Herbert S. Wilf. An open formula for the numerical integration of first order differential equations (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 11(59):201– 203, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wilson:1957:CWW

[Wil57b] W. L. Wilson, Jr. Comment: W. L.

Wilson, Jr., Tables of Inverses to Laplacian Operators over Triangular Grids, two typewritten leaves 81 × 11 inches and one transparency 27 × 48 inches deposited in UMT File. Mathematical Tables and Other Aids to Computation, 11(58):108, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http:

//www.jstor.org/stable/2002157.

Wilf:1958:OFN

[Wil58] Herbert S. Wilf. An open formula for the numerical integration of first order differential equations (II). Mathematical Tables and Other Aids to Computation, 12(61): 55–58, January 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wolf:1952:ONR

[Wol52] J. Jay Wolf. The Office of Naval Research relay computer. Mathematical Tables and Other Aids to Computation, 6(40):207–212, October 1952. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Worsley:1955:SND

[Wor55] Beatrice H. Worsley. Solutions of a nonlinear differential equation arising in the theory of diffusion flames. Mathematical Tables and Other Aids to Computation, 9(51): 112–116, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wilkes:1950:EED

[WR50] M. V. Wilkes and W. Renwick. The EDSAC (electronic delay storage automatic calculator). Mathematical Tables and Other Aids to Computation, 4(30):61–65, April 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wilkes:1957:NMC

[WW57] M. V. Wilkes and D. J. Wheeler. Note on “A method for computing certain inverse functions” (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 11(59):204, July 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wynn:1956:DCT

[Wyn56] P. Wynn. On a device for computing the e_{m}(S_{n}) transformation. Mathematical Tables and Other Aids to Computation, 10(54):91–96, April 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Zondek:1959:EPA

[ZS59] B. Zondek and J. W. Sheldon. On the error propagation in Adams’ extrapolation method. Mathematical Tables and Other Aids to Computation, 13(65):52–55, January 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).